### v1

# Prioritized-MVBA: A New Approach to Design an Optimal Asynchronous Byzantine Agreement Protocol

Abstract
The multi-valued byzantine agreement protocol (MVBA) in the authenticated setting has been widely used as a core to design atomic broadcast and fault-tolerant state machine replication protocols in asynchronous networks. Originating from the seminal work of Cachin et al. [13], subsequent research endeavors have sought to optimize protocol efficiency in terms of communication complexity. Notable advancements following Cachin’s contributions include: i) VABA [23], requiring multiple protocol instances to achieve agreement on a party’s request, and ii) Dumbo-MVBA [29], employing a cryptographic asynchronous dispersal and recovery methods to manage communication complexity alongside additional computational and communication rounds overheads.
Our objective is to devise an MVBA protocol that achieves agreement in each instance without extra computation and communication rounds while maintaining the optimal metrics. Through meticulous research, we analyze the network/adversary’s message distribution patterns and observe how a party’s message reception varies with the number of request senders. We ascertain that opting for a subset of parties ($f+1$, ensuring the inclusion of at least one honest party, at most $f$ number of parties can be faulty ) rather than the entire set ($n$) can facilitate the design of a protocol meeting the aforementioned criteria.
Central to our design approach is the introduction of the committee in the classic MVBA protocol, wherein a randomly selected subset of ($f+1$, where $n=3f+1$) parties get selected and simultaneously broadcast their requests (transactions) to gather verifiable proofs. Successive distributions of these proofs afford us the necessary properties to employ the asynchronous binary Byzantine agreement (ABBA) protocol for reaching an agreement on a selected party’s requests. By integrating the committee and ABBA protocols, we devise the optimal MVBA protocol, termed pMVBA (Prioritized-MVBA). This protocol exhibits resilience to tolerate up to $\lfloor\frac{n}{3}\rfloor$ Byzantine failures, with an expected runtime of $O(1)$, optimal message complexity of $O(n^{2})$, and optimal communication complexity $O((l+\lambda)n^{2})$ (where $n$ represents the number of parties, $l$ denotes the bit length of the input, and $\lambda$ signifies the security parameter).
Keywords Distributed Systems  $\cdot$ Blockchain  $\cdot$ Consensus Protocols
1 Introduction
Byzantine Agreement (BA) protocols are fundamental to implementing decentralized infrastructures, particularly in the context of blockchain and decentralized applications. These protocols ensure agreement among parties, even in the presence of malicious actors, making them crucial for maintaining the integrity and functionality of distributed systems. The renewed interest in BA protocols has been driven by the success of Bitcoin [35] and other decentralized applications [44], which rely heavily on these protocols for achieving agreement [14]. Despite the existence of practical BA protocols [1, 33, 5, 37, 41, 31], their reliance on time parameters for ensuring liveness poses challenges in designing agreement protocols. The FLP impossibility result [32] demonstrates that deterministic BA protocols cannot guarantee agreement in asynchronous networks. This necessitates the development of efficient asynchronous Byzantine Agreement protocols [23, 42, 11, 5, 29, 10, 39, 45, 22] that can overcome the limitations of existing approaches, particularly in terms of communication complexity.
Achieving agreement in asynchronous networks efficiently remains a significant challenge. Classic Multi-Valued Byzantine Agreement (MVBA) protocol like Cachin’s [13], while effective, suffer from high communication complexity, often involving terms like $O(n^{3})$. This high complexity is impractical for large-scale systems, necessitating new approaches to reduce communication overhead while ensuring optimal resilience and correctness. Cachin et al. [13] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision. Cachin’s MVBA protocol provides polynomial-time complexity but suffers from high communication overhead, with a complexity of $O(n^{3})$. VABA [23] introduced a view-based approach to eliminate the $O(n^{3})$ term, while Dumbo-MVBA [29] used erasure codes to handle large input sizes more efficiently, reducing the communication complexity to $O(ln+\lambda n^{2})$. However, these optimizations often introduced additional rounds of communication and computation.
In this context, we propose a novel approach to optimizing MVBA protocols by introducing the Prioritized Multi-Valued Byzantine Agreement (pMVBA) protocol. Our goal is to achieve agreement in each instance without the need for extra computation and communication rounds, maintaining optimal performance metrics. The key innovation of our approach is the integration of a committee-based method within the classical MVBA framework, where a randomly selected subset of parties, rather than the entire set of parties, broadcast proposals.
The pMVBA protocol leverages the concept of a committee, wherein a subset of $f+1$ parties is chosen to broadcast their proposals. This approach ensures that at least one honest party is always included in the subset, enhancing the protocol’s resilience. By dynamically selecting parties for each protocol instance, we mitigate the risk of adversarial attacks targeting specific nodes. This method significantly reduces the number of broadcast messages and the associated communication overhead. Central to our protocol is the integration of the Asynchronous Binary Byzantine Agreement (ABBA) protocol, which facilitates reaching an agreement on one of the proposals broadcast by the committee members. The ABBA protocol ensures that the selected parties’ proposals are agreed upon with probability $1$, maintaining the integrity and functionality of the distributed systems.
Our pMVBA protocol achieves several key improvements over existing MVBA protocols while exhibiting optimal resilience against Byzantine failures, with an expected runtime of $O(1)$, optimal message complexity of $O(n^{2})$, and optimal communication complexity of $O((l+\lambda)n^{2})$, where $n$ is the number of parties, $l$ is the bit length of the input, and $\lambda$ is the security parameter. The key improvements include removing the need for multiple instances of the protocol or extra rounds of messages and cryptographic computation to reach an agreement on a party’s request while keeping the optimal communication complexity. These enhancements make the pMVBA protocol suitable for large-scale decentralized applications, addressing the scalability challenges faced by traditional MVBA protocols.
The main contributions of this paper are as follows:
  • We introduce a novel committee selection protocol that ensures the presence of at least one honest party in the selected subset, enhancing the protocol’s resilience against Byzantine failures.
  • We propose the pMVBA protocol, which reduces the communication complexity from $O(n^{3})$ to $O((l+\lambda)n^{2})$ by dynamically selecting a subset of parties to broadcast their proposals.
  • Our protocol integrates the ABBA protocol to achieve agreement with probability $1$ on the proposals broadcast by committee members, ensuring robust agreement mechanisms.
  • We provide a thorough theoretical analysis demonstrating that the pMVBA protocol achieves an expected constant asynchronous round count while maintaining optimal resilience and correctness.
  • We present extensive case studies and performance evaluations, showing that the pMVBA protocol is effective and scalable for large-scale decentralized applications, outperforming traditional MVBA protocols in terms of communication overhead and efficiency.
2 Related Work
This section provides an overview of the Byzantine Agreement protocols and the key developments in the Byzantine Agreement protocols, focusing on four main areas: Byzantine Agreement Protocols in partially-synchronous network, MVBA Protocols in asynchronous network, Committee-Based Protocols, and Optimistic Fastlane Mechanisms to combine both the partially-synchronous network and the asynchronous network.
Byzantine Agreement Protocols. Byzantine Agreement (BA) protocols have long been a cornerstone of fault-tolerant distributed computing. The Byzantine Generals Problem, introduced by Lamport et al. [26], laid the theoretical foundation for BA protocols, emphasizing the challenges of achieving an agreement in the presence of faulty or malicious nodes. Early BA protocols, such as those by Pease et al. [38], focused on synchronous networks where communication occurs in fixed rounds. However, these protocols often relied on time assumptions for ensuring liveness.
Practical Byzantine Fault Tolerance (PBFT), proposed by Castro and Liskov [15], was a significant advancement in making BA protocols practical for real-world applications. PBFT introduced an efficient consensus protocol for partially-synchronous networks, ensuring safety and liveness even with up to $f<\frac{n}{3}$ Byzantine faults. Several modern BA protocols build upon the principles of PBFT, including HotStuff [46] and Tendermint [12], optimizing various aspects such as communication complexity and responsiveness.
The FLP impossibility result [18] demonstrated that no deterministic BA protocol could guarantee an agreement in an asynchronous network with a single faulty node. This led to the exploration of probabilistic and randomized approaches for ABA. Ben-Or’s protocol [8] was one of the earliest randomized BA protocols, using coin-flipping to break symmetry among nodes. Recent protocols like HoneyBadgerBFT [34] and Dumbo [28] have focused on optimizing ABA for high throughput and low latency.
Multi-Valued Byzantine Agreement Protocols. Cachin et al. [13] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision. Cachin’s MVBA protocol provides polynomial-time complexity but suffers from high communication overhead, with a complexity of $O(ln^{2}+\lambda n^{2}+n^{3})$.
Subsequent works, such as VABA [23] and Dumbo-MVBA [29], aimed to reduce the communication complexity of MVBA. VABA introduced a view-based approach to eliminate the $O(n^{3})$ term, while Dumbo-MVBA used erasure codes to handle large input sizes more efficiently, reducing the communication complexity to $O(ln+\lambda n^{2})$. However, these optimizations often introduced additional rounds of communication and expensive cryptograhic computations.
Our focus on optimizing the classical MVBA protocol aims to address these challenges by reducing the number of broadcasting parties, thus lowering communication overhead while maintaining correctness and resilience.
Committee-Based Protocols. Committee-based protocols have emerged as a promising approach to reduce communication complexity in BA protocols. Algorand [20] introduces the concept of committee members, wherein a predefined set of parties collaborates to reach an agreement. Algorand operates within a permission-less (non-trusted) setting. In a similar vein, COINcidence [19] presents a protocol within a trusted setup, leveraging committee-based participation to achieve agreement with high probability.
Our approach diverges from both Algorand and COINcidence. Unlike Algorand, we operate within a trusted setup, and unlike both Algorand and COINcidence, our protocol involves selected parties undertaking a singular step. Furthermore, our protocol guarantees an agreement on a single party’s requests with a probability of 1. Conversely, FasterDumbo [11] utilizes the committee concept to minimize the number of components in the agreement process, yielding outputs from $n-f$ parties’ requests. Beat [42] amalgamates the most effective elements of HoneyBadgerBFT’s concrete implementation to showcase superior performance across varied settings.
Optimistic Fastlane Mechanisms. Certain works, such as those by Victor et al. [25] and Cachin et al. [17], have explored the incorporation of an optimistic ’fastlane’ mechanism, designed for operation within partially synchronous networks. However, in instances of sluggish agreement processes, these mechanisms transition to asynchronous Byzantine agreement protocols. Recent endeavors, such as those by Joleton [39] and Boltdumbo [45], have adopted a similar strategy. Our work remains pertinent within this context as we rely on asynchronous protocols when the network exhibits Byzantine behavior, particularly within the pessimistic path.
Given our protocol’s focus on outputting requests from a single party, we omit comparisons with DAG-rider [22] and Aleph [21], which utilize directed acyclic graphs for agreement alongside sequential ACS. Additionally, recent efforts have aimed to eliminate the private setup phase in asynchronous Byzantine agreement protocols, emphasizing asynchronous distributed key generation [40, 30, 7].
3 System Model
We use the standard notion [16, 13] to describe a distributed algorithm involving $n$ parties and an adversary in an authenticated setting. This section explains the foundational assumptions, network model, adversary model, computational model, and cryptographic tools used in our protocol.
3.1 System and Network Assumptions
Parties and Their Setup: There are $n$ parties, denoted as $p_{1},p_{2},\ldots,p_{n}$. We consider a trusted setup of threshold cryptosystems. The system provides a secret key and a public key to each party before the protocol starts. A party uses its secret key to sign a message, and other parties use the corresponding public key to verify the signed message. The generation and distribution of these keys are out of the scope of this paper. We follow standard literature for key generation and distribution and refer interested readers to [9, 4, 6]. We use the node and party alternatively throughout the paper.
Network Model: Parties are connected via reliable, authenticated point-to-point channels. Here, reliable implies that if an honest party $p_{i}$ sends a message to another honest party $p_{j}$, the adversary cannot modify or drop the message. However, the adversary can delay the message delivery to influence the protocol execution time. Since we consider an adaptive adversary that can corrupt a party at any time during the protocol execution, the adversary can corrupt a party $p_{i}$ after it sends a message and then make the party $p_{i}$ drop the message.
Adversary Model: We consider an adaptive adversary that can corrupt any party during the protocol execution. If the adversary takes control of a party, the party is corrupted, reveals its internal state to the adversary, behaves arbitrarily, and remains corrupted. An honest party follows the protocol, keeps its internal state secret from the adversary, and remains uncorrupted throughout the protocol execution. The adversary can corrupt $f$ parties among the $n$ parties, where $f<\frac{n}{3}$.
Computational Model: We adopt standard modern cryptographic assumptions and definitions from [16, 13]. The assumptions allow the parties and the adversary to be probabilistic polynomial-time interactive Turing machines. This means that upon receiving a message, a party carries out some computations, changes its state, generates outgoing messages if required, and waits for the next incoming message. To rule out infinite protocol executions and restrict the run time of the adversary, we require the message bits generated by honest parties to be probabilistic uniformly bounded by a polynomial in the security parameter $\lambda$. Therefore, we assume that the number of parties $n$ is bounded by a polynomial in $\lambda$.
3.2 Validated Asynchronous Byzantine Agreement:
Multi-valued Byzantine Agreement (MVBA) allows parties to agree on an arbitrary string $\{0,1\}^{l}$, which must satisfy a predefined validity condition before the honest parties agree on it. This protocol ensures that the agreed-upon values are valid even when inputs come from malicious parties. This property is crucial for ensuring the protocol’s correctness. Therefore, it is called the validated asynchronous byzantine agreement [13, 24, 47].
Multi-valued Byzantine Agreement The MVBA protocol provides a polynomial-time computable predicate $Q$. Each party inputs a value $v$ that must satisfy the condition specified by $Q$, which is application-dependent. The MVBA protocol guarantees the following properties with negligible probability of failure:
  • Liveness. If all honest parties are activated, then all honest parties will reach a decision.
  • External Validity. If an honest party outputs a value $v$, then $v$ satisfies the predicate $Q(v)=\text{true}$.
  • Agreement. If two honest parties output values $v$ and $v^{\prime}$, respectively, then $v=v^{\prime}$.
  • Integrity. If honest parties output a value $v$, then $v$ must have been input by a party.
  • Efficiency. The number of messages generated by honest parties is uniformly bounded with high probability.
3.3 Preliminaries
Asynchronous Binary Byzantine Agreement Biased Towards 1: The ABBA protocol allows parties to agree on a single bit $b\in\{0,1\}$.
The ABBA protocol guarantees the following properties. Additionally, the biased external validity property applies to the biased ABBA protocol.
  • Agreement. If an honest party outputs a bit $b$, then every honest party outputs the same bit $b$.
  • Termination. If all honest parties receive input, then all honest parties will output a bit $b$.
  • Validity. If any honest party outputs a bit $b$, then $b$ was the input of at least one honest party.
  • Biased External Validity. If at least $\langle f+1\rangle$ honest parties propose $1$, then any honest party that terminates will decide on $1$.
(1, $\kappa$, $\epsilon$)- Committee Selection: The committee selection protocol is executed among $n$ parties (identified from 1 through $n$). This protocol ensures that an honest party outputs a $\kappa$-size committee set $C$ with at least one honest member, given that at least $f+1$ honest parties participate. A protocol is a $(1,\kappa,\epsilon)$-Committee Selection protocol if it satisfies the following properties with negligible probability of failure in the cryptographic security parameter $\lambda$:
  • Termination. If $\langle f+1\rangle$ honest parties participate in the committee selection and the adversary delivers the messages, then the honest parties will output the set $C$.
  • Agreement. Any two honest parties will output the same set $C$.
  • Validity. If any honest party outputs the set $C$, then:
    • (i) $|C|=\kappa$,
    • (ii) the probability of every party $p_{i}$ being in $C$ is the same, and
    • (iii) $C$ contains at least one honest party with probability $1-\epsilon$.
  • Unpredictability. The probability of the adversary predicting the committee $C$ before any honest party participates is at most $\frac{1}{{}^{n}C_{\kappa}}$.
Guo et al. [11] constructed the $(1,\kappa,\epsilon)$-Committee Selection protocol using a threshold coin-tossing mechanism (see Appendix A.3), which is derived from threshold signatures. The protocol ensures that at least one honest party is a committee member with overwhelming probability $1-\epsilon-\text{neg}(\lambda)$, where $\text{neg}(\lambda)$ is a negligible function in the cryptographic security parameter $\lambda$, and $\epsilon$ is $exp(-\Omega\kappa)$.
Cryptographic Abstractions: We design a distributed algorithm in authenticated settings where we use robust, non-interactive threshold signatures to authenticate messages and a threshold coin-tossing protocol to select parties randomly [16].
  1. 1.
    Threshold Signature Scheme: We utilize a threshold signature scheme introduced in [43, 16]. The basic idea is that there are $n$ parties, up to $f$ of which may be corrupted. The parties hold shares of the secret key of a signature scheme and may generate shares of signatures on individual messages. $t$ signature shares are both necessary and sufficient to construct a signature where $f<t\leq(n-f)$. The threshold signature scheme also provides a public key $pk$ along with secret key shares $sk_{1},\ldots,sk_{n}$, a global verification key $vk$, and local verification keys $vk_{1},\ldots,vk_{n}$. Initially, a party $p_{i}$ has information on the public key $vk$, a secret key $sk_{i}$, and the verification keys for all the parties’ secret keys. We describe the security properties of the scheme and related algorithms in Appendix A.2.
  2. 2.
    Threshold Coin-Tossing Scheme: The threshold coin-tossing scheme, as introduced in [43, 16], involves parties holding shares of a pseudorandom function $F$ that maps the name $C$ (an arbitrary bit string) of a coin. We use a distributed pseudorandom function as a coin that produces $k^{\prime\prime}$ random bits simultaneously. The name $C$ is necessary and sufficient to construct the value $F(C)\in\{0,1\}^{k^{\prime\prime}}$ of the particular coin. The parties may generate shares of a coin — $t$ coin shares are both necessary and sufficient where $f<t\leq n-f$, similar to threshold signatures. The generation and verification of coin-shares are also non-interactive. We describe the security properties of the scheme and related algorithms in Appendix A.3.
4 The Proposed Protocol
(a) An overview of the proposed protocol framework, highlighting our contributions: committee selection, the VCBC protocol for committee members, recommendations, and Sequential ABBA.
In this section, we present a protocol for multi-valued Byzantine agreement capable of tolerating up to $f<\frac{n}{3}$ Byzantine faults. The protocol features an expected communication bit complexity of $O(ln^{2}+\lambda n^{2})$ and an expected constant asynchronous round count. Our modular implementation comprises five distinct sub-protocols: committee selection, prioritized verifiable consistent broadcast protocol (pVCBC), Propose-Recommend, random order/permutation generation, and Sequential-ABBA. The framework of the proposed protocol is illustrated in Figure 1(a). The violet-colored boxes indicate our contributions to the protocol. We break down the protocol into two main parts. The top part includes the committee selection and the VCBC protocol for the committee members, while the bottom part handles the distribution of the committee members’ messages and the agreement protocol. Section 4.1 provides an overview and the construction of party selection and broadcast (top part of the framework), and Section 4.2 offers a detailed exposition of the protocol, including pseudocode and descriptions.
4.1 Optimizing Message and Communication Overhead
To enhance the efficiency of Byzantine agreement protocols, we propose a novel method to minimize the number of broadcast requests. Rather than requiring all $n$ parties to broadcast their requests, we select a subset of parties to do so, ensuring that the protocol’s progress and security properties are maintained.
Party Selection and Broadcast: In traditional asynchronous Byzantine agreement protocols, all $n$ parties are typically allowed to broadcast their proposals. However, the main factor contributing to the $O(n^{3})$ communication complexity in classic MVBA protocols is that each of the $n$ parties must broadcast an $O(n)$-size array. These arrays include broadcast information from all parties. Our approach aims to reduce the number of broadcasting parties, thereby eliminating the need for transmitting an $O(n)$-size array and decreasing the number of required Asynchronous Binary Byzantine Agreement (ABBA) instances. From this point forward, we use proposal instead of request and propose instead of broadcasting requests.
To achieve this reduction, we select at least $f+1$ parties to broadcast their proposals, ensuring that there is at least one honest party among them. This selection is crucial for guaranteeing the protocol’s progress. We employ a committee selection protocol similar to FasterDumbo [11], which uses a cryptographic coin-tossing scheme to randomly select $\kappa$ (a security parameter) parties. We set $\kappa=f+1$ to ensure the inclusion of at least one honest party, increasing the likelihood of multiple honest parties being selected.
Dynamic selection of parties for each protocol instance prevents adversaries from corrupting fixed parties, ensures fair opportunity for all parties to propose, and mitigates the risk of starvation and Denial-of-Service (DoS) attacks. The committee selection process must be immune to adversary influence. The cryptographic coin-tossing scheme ensures random selection, requiring honest party participation after the completion of the previous instance. We used the committee member/selected party alternatively.
(a) Committee-Selection protocol. Each party shares their $s_{i}=CShare(r_{i})$ and a party requires f+1 shares. We see a party use $CToss(r_{1},s)$ function to get the selected parties and each party gets the same set of parties. We show only for the two parties but it is true for every party.
(b) pVCBC protocol. Each selected party (here $p_{1}$ and $p_{2}$ ) propose its requests. Each receiving party verify the requests and add its sign-share on the message and returns to the sender. A party waits for 2f+1 sign-shares (here it is 3 because f = 1). The 2f+1 sign shares are necessary to create a threshold-signature ($ts1$).
Committee Selection Protocol: The committee selection protocol (CS) is an integral part of our approach. It ensures that a reliable subset of parties is chosen to broadcast proposals, thus reducing communication complexity. The steps of the CS protocol are illustrated in Algorithm 1 and described as follows:
  • Generating Coin-Shares: Upon invocation, a party generates a coin-share $\sigma_{i}$ for the current instance and broadcasts it to all parties, waiting to receive $f+1$ coin-shares. (lines 3-5)
  • Verifying Coin-Shares: When a party receives a coin-share for the first time from a party $p_{k}$, it verifies the coin-share and accumulates it in set $\Sigma$. This process continues until $f+1$ valid shares are collected. (lines 8-10)
  • Selecting Parties: After receiving $f+1$ valid coin-shares, a party uses the $CToss$ function along with the collected coin-shares to select $f+1$ parties. (lines 6-7)
1
2 Local variables initialization:
3      $\Sigma\leftarrow\{\}$
4
5
6      upon $CS\langle id,instance\rangle$ invocation do
7           $\sigma_{i}$leftarrow $CShare\langle r_{i}\rangle$
8           $multi-cast\langle SHARE,id,\sigma_{id},instance\rangle$
9           wait until $|\Sigma|=f+1$
10           return $CToss\langle r_{i},\Sigma\rangle$
11
12
13          upon receiving $\langle SHARE,id,\sigma_{k},instance\rangle$ from a party $p_{k}$ for the first time do
14                if $CShareVerify\langle r_{k},k,\sigma_{k}\rangle=true$  then
15                     $\Sigma\leftarrow{\sigma_{k}\cup\Sigma}$
16
17
Algorithm 1 Committee Selection Protocol for party $p_{i}$
We provide a illustration of the CS protocol in figure 2(a) and the cryptographic coin-tossing definitions (e.g. $CShareVerify$) in the Appendix A.3.
Broadcast protocol A committee member requires a verifiable proof of its proposal to demonstrate that it has broadcast the same proposal to at least $f+1$ honest parties. Typically, the VCBC is employed for this purpose, allowing every party to generate a verifiable proof. Since our protocol restricts proposal broadcasting to only the selected parties, we employ a slightly modified version of the verifiable consistent broadcast protocol, termed pVCBC. This adaptation ensures that if a party receives a verifiable proof from another party, it does not need to verify whether the proof originates from a selected party or not. The definition of the VCBC protocol and the construction of the pVCBC protocol can be found in the Appendix A.1 and Appendix C. We have also provided an illustration of the protocol in Figure 2(b).
4.2 Prioritized MVBA
This section details the MVBA module in an asynchronous network. We propose a novel MVBA protocol called Prioritized-MVBA (pMVBA), which uses a committee as a prioritized parties and reduces the communication complexity of the classic MVBA protocol.
Figure 3: New-VBAConst protocol. After completing the pVCBC broadcast, each party proposes its value along with verifiable proof. Upon receiving verifiable proof, the party recommends the received proposal. Each party waits for $n-f$ recommendations, then uses the coin-tossing scheme (CShare) to generate a permutation and initiates the ABBA protocol to reach an agreement on one of the proposals.
Overview of pMVBA The pMVBA protocol is inspired by the asynchronous binary Byzantine agreement methodology from Cachin et al. [13]. It significantly reduces the number of initial broadcasts, eliminating costly sub-components by leveraging the message distribution pattern of the adversary.
In classic MVBA, each party broadcasts an $O(n)$-size array to achieve $O(1)$ expected constant runtime. Our pMVBA protocol as shown in Fig 3 and explained in Section 4.1, however, restricts this broadcasting to $f+1$ parties, significantly reducing communication complexity. The protocol replaces the $O(n)$ commit broadcast with a single recommend message. We provide a details about classic MVBA protocol challenges in the Appendix B.2.
Since $f+1$ parties generate verifiable proofs and broadcast their proposals, a party must respond after receiving the first verifiable proof. If multiple parties complete the pVCBC protocol, different parties may receive proofs from different parties, hindering agreement. Therefore, we introduce the recommend step, where a party broadcasts received proofs to all parties, ensuring at least one proof reaches the majority of the parties ($12f+$ 1), which is necessary to employ ABBA protocol to reach an agreement. We use majority and threshold-number alternatively.
Replacing Commit with Recommend The key idea in pMVBA is to replace the commit step with the recommend step. The commit step in classic MVBA protocols involves collecting $n-f$ verifiable proofs and creating an array of length $n$. In contrast, the recommend step involves broadcasting verifiable proofs directly, significantly reducing complexity.
Construction of the Recommend Step The recommend step consists of sending and collecting recommendations. This step ensures that at least one verifiable proof reaches the majority of parties. The detailed SCR protocol is shown in Algorithm 2:
  • Initialization: A party initializes an empty set $\Sigma$.
  • Broadcast Recommendation: A party creates a recommendation message and broadcasts it to all parties.
  • Recommendation Collection: Upon receiving a recommendation, a party verifies the threshold-sign and instance. If valid, the party adds the proposal and $\sigma$ to its set $\Sigma$.
  • A party waits for $n-f$ recommendation messages before proceeding.
1 Local variables initialization:
2 $\Sigma\leftarrow\{\}$
3
4 upon $SCR\langle ID,proposal,\sigma\rangle$ invocation do
[…]
6       wait until $|\Sigma|$ = $n-f$
7       return $\Sigma$
8
9
10
11upon receiving $\langle recommendation,ID,proposal,\sigma\rangle$ from party $p_{j}$ for the first time do
12       if  $ID.instance=instance$wedge $Verify_{id}\langle proposal,\sigma\rangle=true$  then
13             $\Sigma\leftarrow{\langle proposal,\sigma\rangle\cup\Sigma}$
14
15
16
Algorithm 2 Recommend Protocol for all parties
Detailed pMVBA Protocol We now present the full details of the pMVBA protocol, inspired by the Cachin-MVBA protocol [13] but with fewer broadcasts. The protocol replaces the $O(n)$ commit broadcast with one recommend message and runs an expected constant number of ABBA instances to reach an agreement on a party’s proposal.
The formal description of the pMVBA protocol is shown in Algorithm 3. It consists of the following steps:
  • Party/Committee Selection: When a new instance starts, a party invokes the CS protocol to select parties. (See Algorithm 1 lines 1-10 and Algorithm 3 line 10)
  • pVCBC Broadcast: A selected party broadcasts its proposal using the pVCBC protocol and generates verifiable proofs of their broadcasts. (See Algorithm 3 line 13 and Algorithm 4 lines 1-16.)
  • Propose and Recommend: Upon completing the pVCBC protocol, a party broadcasts (as a propose) its verifiable proof. When a party receives a verifiable proof, it broadcasts the proof as a recommendation and a party waits for $n-f$ recommendation messages. (See Algorithm 3 lines 14-22 and the Algorithm 2)
  • Permutation: When a party receives $n-f$ recommendations, it generates a random order of the selected parties using the cryptographic coin-tossing scheme. (See Algorithm 3 line 22 and Algorithm 4 lines 19-29 )
  • Sequential-ABBA: For the agreement phase, a party runs the ABBA protocol for the selected parties in the determined order, stopping once an agreement is reached. (See Algorithm 3 lines 32-58 and Algorithm 5 lines 1-28).
1
2 Local variables initialization:
3      $instance\leftarrow 1$
4      $PParties\leftarrow\bot$
5      $\sigma\leftarrow\bot$
6      $proposal\leftarrow\bot$
7      $W\leftarrow\bot$
8      $\Sigma\leftarrow\bot$
9
10
11      while true do
12           $ID\leftarrow{\langle instance,id\rangle}$
13           $PParties\leftarrow CS\langle ID,instance\rangle$
14
15          if $p_{i}\in PParties$ then
16                $proposal\leftarrow request$
17                $\sigma\leftarrow pVCBC\langle ID,requests\rangle$
[…]
19
20                wait for a propose or a recommend message.
21               upon receiving $\langle\textit{propose},ID,\sigma,proposal\rangle$ for the first time do
22                     $W\leftarrow SCR\langle ID,proposal,\sigma\rangle$
23
24
25                    upon receiving $\langle\textit{recommendation},ID,proposal,\sigma\rangle$ for the first time do
26                          if no $recommendation$ is sent yet  then
27                               $W\leftarrow SCR\langle ID,proposal,\sigma\rangle$
28
29
30
31                              Permutation:
32                               $\Sigma\leftarrow Permutation\langle ID,PParties\rangle$
33
34                              Sequential-ABBA:
35                               $proposal\leftarrow Sequential-ABBA\langle\Sigma,W\rangle$
36                               $decide\langle proposal\rangle$
37                               $instance\leftarrow instance+1$
38
39
Algorithm 3 Prioritized-MVBA: Protocol for party $p_{i}$
5 Evaluation and Analysis of the Proposed Protocol
This section evaluates the proposed protocol based on its correctness and efficiency. Correctness ensures that the protocol maintains predefined security properties, making it resilient against adversarial attacks. Efficiency evaluates the protocol’s performance in terms of resource utilization and execution time. In this Section, we conduct extensive evaluation and analysis to answer:
•  What are the formal proofs supporting the correctness of the Prioritized-MVBA protocol? (§$\S$ 5.1)
•  How does the Prioritized-MVBA protocol perform in terms of message complexity, communication complexity, and running time? (§$\S$ 5.2)
•  What are the observed behaviors and outcomes of the protocol under different network conditions and adversarial strategies? (§$\S$ 5.3)
•  How does the performance of the Prioritized-MVBA protocol compare to other MVBA protocols in terms of key metrics? (§$\S$ 5.4)
5.1 Proof of the Prioritized-MVBA Protocol
The correctness of the proposed protocol is critical to ensure that it adheres to the standard MVBA outcomes. Our goal is to prove that reducing the number of broadcasts does not compromise the protocol’s reliability. The output of a party’s request depends on the output of an ABBA instance within our protocol’s agreement loop, typically resulting in a value of 1. For the ABBA protocol to output 1, at least one honest party must input 1. We substantiate our protocol’s integrity by demonstrating that at least $12f+$ 1 parties receive a verifiable proof of a party’s proposal, thus ensuring the necessary input for the ABBA instance.
Lemma 5.1.
At least one party’s proposal reaches $12f+$ 1 parties.
Proof.
Since the selected parties can be byzantine and the adversary can schedule the message delivery to delay the agreement, we have considered the following scenarios:
  1. 1.
    Among $\langle f+1\rangle$ selected parties, $f$ parties are non-responsive.
  2. 2.
    Selected $\langle f+1\rangle$ parties are responsive, but other $f$ non-selected parties are non-responsive.
  3. 3.
    Every party is responsive, including the selected $\langle f+1\rangle$ parties.
  4. 4.
    Selected $t<\langle f+1\rangle$ parties are responsive, and total $m$ parties responsive, where $\langle 2f+1\rangle\leq m\leq n$.
We prove that in every scenario, at least one party’s proposal reaches $\langle 2f+1\rangle$ parties.
  1. 1.
    Since among $\langle f+1\rangle$ selected parties, $f$ parties are non-responsive, only one party completes the $pVCBC$ protocol and broadcasts the verifiable proof. If any party receives a verifiable proof, then the verifiable proof is from the responsive selected party. Since every party receives the verifiable proof for the same party’s proposal, the proposal reaches $\langle 2f+1\rangle$ parties.
  2. 2.
    The $\langle f+1\rangle$ selected parties are responsive and complete the $pVCBC$ protocol. Each selected party broadcasts the verifiable proof, and $\langle 2f+1\rangle$ parties receive the verifiable proof ( another $f$ number of parties are non-responsive). Any party receives a verifiable proof, recommends the received verifiable proof, and waits for $\langle 2f+1\rangle$ recommendations. If a party receives $\langle 2f+1\rangle$ recommendations, then these recommendations include all $\langle f+1\rangle$ parties’ verifiable proofs because among the $\langle 2f+1\rangle$ received recommendations, $\langle f+1\rangle$ number of recommendations are from the selected parties. So, every proposal reaches $\langle 2f+1\rangle$ parties.
  3. 3.
    The proof is by contradiction. Let no proposal reaches more than $2f$ parties. Since we assume every party is responsive, every party receives a proposal in $propose$ step. We also know that no party participate in $Sequential-ABBA$ protocol until the party receives $\langle 2f+1\rangle$ recommendations and get the order to Sequential-ABBA. There must be a $\langle 3f+1\rangle*\langle 2f+1\rangle$ recommendation messages. If no proposal can be recommended to more than $2f$ parties, the total number of recommendations is $\langle 3f+1\rangle*\langle 2f\rangle<\langle 2f+1\rangle*\langle 2f+1\rangle$ (Though a proposal can be recommended by more than one party we assume that every party recommends to the same $2f$ parties otherwise it would fulfil the requirement of $\langle 2f+1\rangle$ proposals). But, the honest parties must send enough recommendation messages to ensure the protocol’s progress and the adversary delivers the messages eventually. Therefore, at least one party’s proposal reaches $\langle 2f+1\rangle$ parties, a contradiction.
  4. 4.
    The proof is by contradiction. Let no proposal reaches to more than $2f$ parties. If $1\leq t\leq\langle f+1\rangle$ parties distribute their verifiable proof to $\langle 2f+1\rangle\leq m<\langle 3f+1\rangle$ parties and no proposal reaches more than $2f$ parties, then there must be no more than $m*2f$ recommendations. However, $m$ parties must receive $m$*$\langle 2f+1\rangle$ recommendations greater than $m*2f$, a contradiction.
This lemma ensures that if $12f+$ 1 parties receive a verifiable proof, at least one honest party will input 1 to the ABBA instance, ensuring the protocol reaches an agreement on 1.
Lemma 5.2.
Without any permutation, the adversary can cause at most $f+1$ iterations of the agreement loop in the Sequential-ABBA protocol.
Proof.
The proof works by counting the total number $A$ of $vote=0$ messages (line 44 of Algorithm 4) that are generated by honest parties (over all 00 vote in the agreement loop). Since every honest party has received a verifiable proof from one party in the $propose$ broadcast step (line 18 of Algorithm 3), the party generates $vote=0$ messages for at most $f$ number of selected parties. Thus $A\leq(f)(n-f)$ (all $(n-f)$ votes are zero). For the asynchronous binary byzantine agreement protocol to decide 00 for a particular party’s proposal and to cause one more iteration of the loop, at least $\langle f+1\rangle$ honest parties must propose 00 for the binary agreement (line 46 of Algorithm 4). Since honest parties only propose 00 if they have received $\langle 2f+1\rangle$ $vote=0$ messages, there must be at least $\langle f+1\rangle$ honest parties who have generated a $vote=0$ message in this iteration. Let $R$ denote the number of iterations of the loop where the binary agreement protocol decides 00. From the preceding argument, we have $A\geq R(f+1)$. Combining these two bounds on $A$, we obtain $R(f+1)\leq f*(n-f)$.
$\begin{split}R\leq\frac{(f)(n-f)}{f+1}\\ R\leq\frac{(f)(2f+1)}{f+1}\\ R\leq\frac{(f)(f+1+f)}{f+1}\\ R\leq f+\frac{f^{2}}{f+1}\newline \end{split}$ (1)
However, there are only $\langle f+1\rangle$ number of selected parties’ proposals, and from Lemma 5.1, at least one verifiable proof reaches $\langle f+1\rangle$ honest parties. Therefore, $R<\langle f+1\rangle$ and the adversary can cause at most $R+1$ or $\langle f+1\rangle$ number of iterations. ∎
Lemma 5.3.
Let $\overline{A}\subseteq\{1,2,...,f+1\}$ be the set of selected parties for which at least $f+1$ honest parties receive the verifiable proof, and let $\Pi$ be a random permutation of the $f+1$ selected parties. Then, except with negligible probability:
  • For every party $p\in\overline{A}$, the ABBA protocol on $ID|p$ will decide $1$.
  • $|\overline{A}|\geq 1$.
  • There exists a constant $\beta>1$ such that for all $t\geq 1$, $Pr[\Pi[1]\notin\overline{A}\wedge\Pi[2]\notin\overline{A}\wedge...\wedge\Pi[t]\notin\overline{A}]\leq\beta^{-t}$.
Proof.
The binary agreement protocol biased towards $1$ decides 00 if the honest parties input 00. From the Sequential-ABBA protocol, each party must receive $\langle n-f\rangle$ $vote=0$ messages in order to input 00 to an ABBA instance and to decide 00. However, from Lemma 5.1 at least $\langle f+1\rangle$ honest parties receive at least one verifiable proof from a selected party $p$, then it is not the case for the party $p\in\overline{A}$. This proves the first claim.
To prove the second claim, let $A$ denote the total number of 00-votes cast by honest parties. Since an honest party votes 00 for at most $f$ number of parties proposals, we have $A\leq f(n-f)$. On the other hand, a party can cast $vote=0$ for $f+1-|\overline{A}|$ parties proposals, and at least $\langle f+1\rangle$ honest parties cast $vote=0$ to decide 00 then $A\geq(f+1$ $|\overline{A}|)(f+1)$. These bounds on $A$ are the same as in Lemma 5.2 with $R=f+1$-$|\overline{A}|$. Using the same argument, it follows that $|\overline{A}|\geq 1$, since $R$ is at most $f$ [^1] To prove the Lemma 5.2 and Lemma 5.3, we have used the same technique of Lemma 7 and Lemma 9 from [13]. We also refer interested readers to Lemma 10 of the same paper for the pseudorandom generation..
The third claim follows now because $|\overline{A}|$ is at least a constant fraction of $f+1$ and thus, there is a constant $\beta>1$ such that $Pr[\Pi(i))\notin\overline{A}]\leq\frac{1}{\beta}$ for all $1\leq i\leq t$. Since the probability of the $t$ first elements of $\Pi$ jointly satisfying the condition is no larger than for $t$ independently and uniformly chosen values, we obtain $Pr[\Pi[1]\notin\overline{A}\wedge\Pi[2]\notin\overline{A}\wedge....\wedge\Pi[t]\notin\overline{A}]\leq\beta^{-t}$.
Theorem 5.4.
Given a protocol for biased binary Byzantine agreement and a protocol for verifiable consistent broadcast, the Prioritized-MVBA protocol provides multi-valued validated Byzantine agreement for $n>3f$ and invokes a constant expected number of binary Byzantine agreement protocols.
Proof.
Agreement: If an honest party outputs a value $v$, then every honest party outputs the value $v$. From Lemma 5.1, at least one proposal reaches $12f+$ 1 parties, and the ABBA protocol reaches an agreement on $1$ if $f+1$ honest parties input $1$. The agreement property of the ABBA protocol ensures all honest parties output $1$ for that ABBA instance and receive the same value $v$, which satisfies $threshold-validate\langle v,\sigma\rangle=true$.
Liveness: If honest parties participate and deliver messages, all honest parties decide. From Lemma 5.1, at least one proposal reaches $12f+$ 1 parties, ensuring the ABBA protocol decides $1$. Lemma 5.3 confirms the protocol reaches an agreement after a constant expected number of iterations.
External-validity: If an honest party terminates and decides on a value $v$, then $externally-valid\langle v,\sigma\rangle=true$. The validity of the ABBA protocol ensures at least one honest party inputs $1$ to the ABBA instance, meaning the honest party received a valid $threshold-signature$ $\sigma$ for $v$.
Integrity: If honest parties decide on a value $v$, then $v$ was proposed by a party. The ABBA protocol returns $1$ if at least one honest party inputs $1$, which requires a valid $threshold-signature$ $\sigma$. Honest parties reply with $sign-shares$ only if a value $v$ is proposed, ensuring $v$ was proposed by a party. ∎
5.2 Efficiency Analysis
The efficiency of a Byzantine Agreement (BA) protocol depends on message complexity, communication complexity, and running time. We analyze the proposed protocol’s efficiency by examining its sub-components: the pVCBC sub-protocol, committee-selection and permutation generation, propose-recommend steps, and the Sequential-ABBA sub-protocol.
Running Time:
Each sub-protocol and step, except for Sequential-ABBA, has a constant running time. The running time of the proposed protocol is dominated by the Sequential-ABBA sub-protocol, which has an expected constant number of asynchronous rounds.
Message Complexity:
In all sub-protocols and steps, except for pVCBC and propose steps, each party communicates with all other parties. Thus, the message complexity is $O(n^{2})$. The expected message complexity of the Sequential-ABBA protocol is also $O(n^{2})$.
Communication Complexity:
The communication complexity of each sub-protocol and step is $O(n^{2}(l+\lambda))$, where $l$ is the bit length of input values and $\lambda$ is the bit length of the security parameter. The expected communication complexity of the Sequential-ABBA protocol is also $O(n^{2}(l+\lambda))$.
In conclusion, the Prioritized-MVBA protocol achieves optimal efficiency and correctness by reducing the number of broadcasts and ensuring all security properties of a multi-valued Byzantine agreement protocol. The detailed analysis confirms its resilience, integrity, liveness, and external validity, making it suitable for robust decentralized infrastructures.
5.3 Case Study
The key contribution of this paper is the reduction in the number of requesting parties and the subsequent utilization of fewer broadcasts to eliminate the expensive sub-components of the classic MVBA protocol. The main challenge is to provide sufficient information to the parties to maintain the protocol’s progress while removing the costly elements. It is crucial to ensure that at least one selected party’s proposal reaches at least $12f+$ 1 parties. This is achieved through the pVCBC step and the propose-recommend step. To illustrate the protocol’s effectiveness, we present a case study in two parts. Section 5.3.1 demonstrates how the protocol achieves the desired properties with the minimum number of nodes a system can have. To further clarify that the protocol can maintain these properties with a larger number of nodes, Section 5.3.2 provides a case study using charts for a system with three times the number of faulty nodes.
5.3.1 Message Flow in Different Stages
In an asynchronous network, the presence of a faulty node and the adversary’s ability to manipulate message delivery can delay the agreement process. This section shows how a faulty party can affect message delivery patterns and how the adversary can delay specific messages to prevent a node’s proposal from reaching the majority. Despite these challenges, the goal is to ensure that nodes can still reach an agreement on a party’s proposal.
Figure 4(a) assumes no faulty nodes, with the adversary delivering messages uniformly, allowing the protocol to reach an agreement on the first try. Figure 4(b) considers a scenario where one node gets a threshold-signature ($ts$) early, or the adversary prioritizes one node’s $ts$ delivery over others, preventing one selected node’s proposal from reaching the majority. Figure 5(a) assumes the selected node $p_{2}$ is either faulty or completely isolated by the adversary, and Figure 5(b) assumes a non-selected node is faulty or isolated, positively impacting the $ts$ delivery of the selected nodes.
(a) Every selected node is non-Byzantine and the adversary delivers the threshold-signature (ts1, ts2) uniformly (both $p_{1}$ and $p_{2}$’s threshold-signature are received by the same number of parties). Both $p_{1}$’s and $p_{2}$’s proposals are received by every party.
(b) Every selected node is non-Byzantine and the adversary delivers the messages non-uniformly (Node $p_{1}$’s threshold-signature is received by three parties but node $p_{2}$’s threshold-signature is received by one party). $p_{1}$’s proposal is received by every party, but node $p_{2}$’s proposal is received by only one party.
(a) One selected node is non-Byzantine ($p_{1}$), and the adversary delivers the messages from this node to every other node. Consequently, only node $p_{1}$ completes the pVCBC protocol and proposes the threshold-signature. Therefore, every party has the proposal except $p_{2}$ (either Byzantine or non-responding). The red node represents either the node is Byzantine or non-responding or the network is not delivering the messages from that node.
(b) Every selected node is non-Byzantine, and the adversary delivers the messages uniformly (both $p_{1}$ and $p_{2}$’s threshold-signature are received by the same number of parties). Both $p_{1}$’s and $p_{2}$’s proposals are received by every party. The non-selected node $p_{4}$ is not responding, which does not affect the overall criteria. The red node represents either the node is Byzantine or non-responding or the network is not delivering the messages from that node.
In conclusion, the above figures collectively demonstrate that regardless of the message delivery pattern or the presence of faulty nodes, the protocol consistently achieves agreement, thereby proving its robustness and effectiveness.
5.3.2 Message Distribution for More Than One Faulty Node
For this study, we assume a total of $n=3f+1=10$ parties, with $f=3$ being faulty. This configuration allows us to explore various message distribution patterns. We require at least one party’s proposal to reach $72f+1=$ 7 parties, with a total of $f+1=4$ selected parties. We examine the impact of different numbers of faulty parties and message distribution patterns. In each figure, the top box indicates selected parties and the number of parties that can recommend the proposal to others (e.g., Party $p_{1}$ can receive 3 recommendations for the first selected party ($R_{1}$), 3 for the second selected party ($R_{2}$), and 2 for the third selected party ($R_{3}$), totaling 7). The bottom box counts the number of recommendations received for each selected party (e.g., $p_{1}$ receives a total of 7 $R_{1}$ recommendations).
Figure 8 assumes no faulty nodes, with the adversary delivering messages uniformly, ensuring that every selected node’s proposal reaches the threshold number of parties. Figure 8 assumes three non-selected non-responding nodes, which also allows every selected node’s proposal to reach the threshold. Figure 8 assumes one honest selected node or the adversary delivering messages from that node, ensuring only the honest selected node’s proposal reaches the threshold. Figure 11 and Figure 11 show scenarios with two and three honest selected nodes, respectively, where their proposals reach the threshold. Figure 11 explores non-uniform message distribution, which may prevent some selected nodes’ proposals from reaching the threshold, but only when all selected nodes are active.
In conclusion, the above scenarios illustrate that the protocol can maintain its effectiveness and achieve agreement even with varying numbers of faulty nodes and different message distribution patterns.
Figure 6: All selected parties complete their pVCBC and broadcast their proposals. All parties are non-Byzantine; therefore, all of them recommend their received proposal. Every party’s proposal reaches at least 7 parties.
Figure 7: All selected parties complete their pVCBC and broadcast their proposals. Three non-selected faulty parties cannot recommend; consequently, a party receives recommendations for all selected parties. Therefore, all selected parties’ proposals reach at least 7 parties.
Figure 8: All selected parties are Byzantine or non-responding except $P_{1}$. The non-Byzantine party completes the pVCBC and proposes its requests. Since there is only one proposal, every party receives the same proposal, which is received by every non-Byzantine party. Therefore, the proposal reaches 7 parties.
Figure 9: Two selected parties are non-Byzantine and complete the pVCBC protocol, proposing their proposals, which are received by every non-Byzantine party. Each selected party’s proposal reaches more than 7 parties.
Figure 10: Three selected parties are non-Byzantine and complete the pVCBC protocol, proposing their proposals, which are received by every non-Byzantine party. Each selected party’s proposal reaches more than 7 parties.
Figure 11: This study tests whether non-uniform message distribution affects recommendation reception. Even with non-uniform distribution, at least half of the selected parties’ proposals reach 7 parties.
5.4 Analysis of Communication Complexity and Resilience
This section evaluates our protocol’s performance by comparing its communication complexity with classic MVBA protocols and assessing its resilience, termination, and safety properties against committee-based protocols.
5.4.1 Comparison of communication complexity with classic protocols.
Our work focuses on the first polynomial time asynchronous byzantine agreement protocol and removes the expensive sub-component from that protcol. We compare our work with the classic MVBA protocol and the traditional MVBA protocols. Table 1 highlights these comparisons. Our techniques differ from the recent improvements in the MVBA protocol. Therefore, we compare our protocol with the classic protocol and the following improvement. Our protocol differs from the Cachin-MVBA [13] protocol in terms of communication complexity. VABA [23] is a view-based protocol where each view is an instance (a complete execution of the protocol) of the protocol. VABA does not guarantee that the parties will reach an agreement at each instance. In our protocol, parties can reach an agreement on a valid proposal in every instance. We also present a comparison between our proposed protocol and the Dumbo-MVBA [29] protocols. The Dumbo-MVBA protocol is intended for large inputs. It uses the erasure code technique to remove the expensive term. The erasure code first disperses the message and then recovers the message at the end of the agreement, which is expensive in terms of latency and computation and is not worth it for small-size inputs. Our protocol achieves the desired communication complexity (removal of the $O(n^{3})$ term) without the utilization of information dispersal techniques (erasure code) and in one instantiation of the protocol. We can use the information dispersal technique like Dumbo-MVBA in our protocol to reduce the communication complexity to $O(ln+\lambda n^{2})$, but our protocol does not depend on the erasure code to remove the $O(n^{3})$ term.
Table 1: Comparison for performance metrics of MVBA protocols
Protocols Comm. (bit) Word Time Msg. ErasureCode. Instance
Cachin [13] $O(ln^{2}+\lambda n^{2}+n^{3})$ $O(n^{3})$ $O(1)$ $O(n^{2})$ No 1
VABA [23] $O(ln^{2}+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ No $\geq 1$
Dumbo-MVBA [29] $O(ln+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ Yes 1
Dumbo-MVBA* [29] $O(ln+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ Yes 1
Our work $O(ln^{2}+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ No 1
5.4.2 Comparison of Resilience, Termination, and Safety with Committee-Based Protocols
We compare our work with notable committee-based protocols, specifically focusing on resilience, termination, and safety properties. Table 2 highlights these comparisons. COINcidence [19] assumes a trusted setup and violates optimal resilience. It also does not guarantee termination and safety with high probability (whp). Algorand [20] assumes an untrusted setup, with resilience dependent on network conditions, and does not guarantee termination whp. The Dumbo [11] protocol uses a committee-based approach, but its committee-election protocol does not guarantee the selection of an honest party, thus failing to ensure agreement or termination. Our protocol achieves optimal resilience and guarantees both termination and safety, as our committee-election process ensures the selection of at least one honest party. This guarantees that the protocol can make progress and reach agreement despite adversarial conditions.
Table 2: Comparison for performance metrics of the committee based protocols
Protocols n> Termination Safety
COINcidence [19] 4.5f whp whp
Algorand [20] * whp w.p. 1
Dumbo1 [11] 3f whp w.p. 1
Dumbo2 [11] 3f whp w.p. 1
Our work 3f w.p. 1 w.p. 1
In conclusion, we have compared our protocol with both the classical protocol and the committee-based protocol. Though our protocol differs from the atomic broadcast protocol in a number of proposals, we provide a comparison of our protocol with the atomic broadcast protocol in Appendix B.1.
6 Discussions and Conclusion
In this paper, we introduced pMVBA, a novel MVBA protocol designed to address the high communication complexity inherent in the classic MVBA protocol. By leveraging a committee-based approach combined with the ABBA protocol, pMVBA significantly reduces communication overhead while maintaining optimal resilience and correctness. Our protocol achieves several key improvements over existing MVBA protocols. By dynamically selecting a subset of parties (f + 1) to broadcast proposals, we ensure that at least one honest party is always included, thereby enhancing the protocol’s resilience. The integration of the pVCBC protocol allows for efficient proposal broadcasting and verifiable proof generation. The recommend step replaces the commit step used in the classic MVBA protocol, reducing the need for extensive communication bit exchanges and achieving agreement with fewer communication rounds. Theoretical analysis and case studies demonstrated the effectiveness of pMVBA, showing that it achieves an expected constant asynchronous round count and optimal message complexity of $O(n^{2})$, where $n$ represents the number of parties. The communication complexity is reduced to $O((l+\lambda)n^{2})$, making pMVBA suitable for large-scale decentralized applications.
Limitations and Future Work. While pMVBA presents significant advancements, it is not without limitations. One limitation is the assumption of a trusted setup for cryptographic keys, which may not be practical in all decentralized environments. Additionally, the protocol’s performance in extremely large networks with highly dynamic membership has not been thoroughly tested and may present scalability challenges. Another limitation is the reliance on the ABBA protocol for agreement, which, while efficient, could still be optimized further to handle more adversarial conditions. Moreover, the security parameters and their impact on the overall performance need more comprehensive analysis under various network conditions. Future work will focus on addressing these limitations. We plan to explore alternative setups that do not require a trusted third party for key distribution, thus enhancing the protocol’s applicability in trustless environments. We will also investigate adaptive mechanisms to improve scalability in large and dynamic networks. Further optimizations to the pMVBA protocol will be considered to enhance its resilience and efficiency. Additionally, extensive empirical evaluations under diverse network conditions will be conducted to better understand the protocol’s performance and robustness.
Appendix A Definitions
A.1 Verifiable Consistent Broadcast
A protocol completes a verifiable consistent broadcast if it satisfies the following properties:
  • Validity. If an honest party sends $m$, then all honest parties eventually delivers $m$.
  • Consistency. If an honest party delivers $m$ and another honest party delivers $m^{\prime}$, then $m=m^{\prime}$.
  • Integrity. Every honest party delivers at most one request. Moreover, if the sender $p_{s}$ is honest, then the request was previously sent by $p_{s}$.
A.2 Threshold Signature Scheme
The $(f+1,n)$ non-interactive threshold signature scheme is a set of algorithms used by $n$ parties, with up to $f$ potentially faulty. The threshold signature scheme satisfies the following security requirements, except with negligible probabilities:
  • Non-forgeability. To output a valid signature, a party requires $t$ signature shares. Therefore, it is computationally infeasible for an adversary to produce a valid signature, as an adversary can corrupt up to $f$ parties ($f<t$) and thus cannot generate enough signature shares to create a valid signature proof for a message.
  • Robustness. It is computationally infeasible for an adversary to produce $t$ valid signature shares such that the output of the share combining algorithm is not a valid signature.
The scheme provides the following algorithms:
[…]
A.3 Threshold Coin-Tossing
We assume a trusted third party has an unpredictable pseudo-random generator (PRG) $G:R\rightarrow\{1,\ldots,n\}^{s}$, known only to the dealer. The generator takes a string $r\in R$ as input and returns a set $\{S_{1},S_{2},\ldots,S_{s}\}$ of size $s$, where $1\leq S_{i}\leq n$. Here, $\{r_{1},r_{2},\ldots,r_{n}\}\in R$ are shares of a pseudorandom function $F$ that maps the coin name $C$. The threshold coin-tossing scheme satisfies the following security requirements, except with negligible probabilities:
  • Pseudorandomness. The probability that an adversary can predict the output of $F(C)$ is $\frac{1}{2}$. The adversary interacts with the honest parties to collect coin-shares and waits for $t$ coin-shares, but to reveal the coin $C$ and the bit $b$, the adversary requires at least $\langle t-f\rangle$ coin-shares from the honest parties. If the adversary predicts a bit $b$, then the probability is $\frac{1}{2}$ that $F(C)=b$ ($F(C)\in\{0,1\}$). Although the description is for single-bit outputs, it can be trivially modified to generate $k$-bit strings by using a $k$-bit hash function to compute the final value.
  • Robustness. It is computationally infeasible for an adversary to produce a coin $C$ and $t$ valid coin-shares of $C$ such that the share-combine function does not output $F(C)$.
The dealer provides a private function $CShare_{i}$ to every party $p_{i}$, and two public functions: $CShareVerify$ and $CToss$. The private function $CShare_{i}$ generates a share $\sigma_{i}$ for the party $p_{i}$. The public function $CShareVerify$ can verify the share. The $CToss$ function returns a unique and pseudorandom set given $f+1$ validated coin shares. The following properties are satisfied except with negligible probability:
  • For each party $i\in\{1,\ldots,n\}$ and for every string $r_{i}$, $CShareVerify(r_{i},i,\sigma_{i})=\text{true}$ if and only if $\sigma_{i}=CShare_{i}(r_{i})$.
  • If $p_{i}$ is honest, then it is impossible for the adversary to compute $CShare_{i}(r)$.
  • For every string $r_{i}$, $CToss(r,\Sigma)$ returns a set if and only if $|\Sigma|\geq f+1$ and each $\sigma\in\Sigma$ and $CShareVerify(r,i,\sigma)=\text{true}$.
Appendix B Miscellaneous
B.1 Comparison with Atomic Broadcast Protocol
As discussed earlier, when the inputs of each party are nearly identical, outputting the requests of $n-f$ parties is not a viable solution. This approach results in higher computational effort without increasing the number of accepted transactions. Table 3 provides a comparison of the communication complexity of our protocol with atomic broadcast protocols. Notably, no atomic broadcast protocol can eliminate the multiplication of $O(n^{3})$ terms. Additionally, atomic broadcast protocols require extra rounds of message exchanges. Here, we focus solely on the communication complexity.
Table 3: Comparison of the communication complexity with the atomic broadcast protocols
Protocols Communication Complexity
HB-BFT/BEAT0 [5] $O(ln^{2}+\lambda n^{3}logn)$
BEAT1/BEAT2 [42] $O(ln^{3}+\lambda n^{3})$
Dumbo1 [11] $O(ln^{2}+\lambda n^{3}logn)$
Dumbo2 [11] $O(ln^{2}+\lambda n^{3}logn)$
Speeding Dumbo [10] $O(ln^{2}+\lambda n^{3}logn)$
Our Work $O(ln^{2}+\lambda n^{2})$
B.2 The Challenge of Classical MVBA Designs
To maintain a message complexity of $O(n^{2})$, the classic MVBA protocol incorporates the Verifiable Consistent Broadcast (VCBC) protocol. Additionally, it introduces the concept of external validity, wherein an input is deemed valid if it satisfies certain criteria. The protocol operates as follows:
Each party utilizes the VCBC protocol to broadcast their request and generate a corresponding verifiable proof. Upon completion of this step, the party broadcasts both the verifiable proof and the request, providing evidence that the request has been broadcast to every other party.
Upon receiving verifiable proof from $n-f$ parties, signaling the completion of the VCBC protocol by the threshold number of parties, a party can initiate the ABBA protocol. However, there exists the possibility that other parties have not received sufficient verifiable proof, or that the adversary manipulates the distribution of proofs in a manner that prevents the majority of ABBA instances from receiving adequate proof.
To address this, each party communicates with others by transmitting an $n$-bit array, indicating receipt of verifiable proof from $n-f$ parties. Upon receipt of $n-f$ verifiable proofs, a party generates a permutation of the parties and invokes ABBA instances based on the order of permutation, ensuring that the number of ABBA instances remains constant on average.
Appendix C Deferred protocols
Construction of the pVCBC
  • Upon invocation of the pVCBC protocol, a party creates a message using the $ID$ and the requests. The party then authenticates the message using its private key and multicasts the message (lines 03-06).
  • Upon receiving a signed message ($msg_{signed}$) from a party $p_{j}$, a party checks whether the sender is a selected party and whether the message is authentic (signed by the sender). If the sender is a selected party and the message is authentic, the party adds its signature share ($sign-share$) to the message, resulting in $\rho$. The party then replies with $\rho$ to the sender (lines 13-16).
  • Upon receiving a signature share $\rho_{k}$ from a party $p_{k}$, a selected party checks the validity of $\rho_{k}$ (ensuring it was added using $p_{k}$’s key). If $\rho_{k}$ is valid, the party adds the signature share $\rho_{k}$ to its set $\Sigma$ (lines 09-11).
  • A selected party waits for $\langle n-f\rangle$ valid signature shares. Upon receiving $\langle n-f\rangle$ signature shares, the party combines them to generate a threshold signature $\sigma$ (proof that the party has sent the same request to at least $\langle f+1\rangle$ honest parties) and returns $\sigma$ to the caller (lines 07-08).
Construction of the Permutation
Here is the pseudocode for the permutation protocol (see Algorithm 4). Below is a step-by-step description of the protocol:
  • Upon invocation of the Permutation protocol, a party generates a coin-share $\sigma_{i}$ for the instance and broadcasts $\sigma_{i}$ to every party, then waits for $12f+$ 1 coin-shares (lines 22-24).
  • When a party receives a coin-share from another party $p_{k}$ for the first time, it verifies the coin-share (ensuring it is from $p_{k}$) and accumulates the coin-share in the set $\Sigma$. The party continues to respond to coin-shares until it has received $12f+$ 1 valid shares (lines 27-29).
  • Upon receiving $12f+$ 1 valid coin-shares, a party uses its $CToss$ function and the collected coin-shares to generate a permutation of the $n$ parties (lines 24-25).
Sequential-ABBA
The purpose of the Sequential-ABBA protocol is to run an agreement loop to agree on a selected party’s proposal. We use the asynchronous binary Byzantine agreement protocol biased towards $1$ for each selected party until the parties reach an agreement on one of the party’s proposals.
Construction of the Sequential-ABBA
Since the Sequential-ABBA protocol guarantees that the parties will eventually reach an agreement on one of the selected party’s proposals, the protocol runs a loop for the selected parties until it agrees on a valid proposal. The pseudocode of the Sequential-ABBA protocol is given in Algorithm 5, and a step-by-step description is provided below:
  • The Sequential-ABBA protocol takes two arguments: (i) $\Sigma$, the permutation of the selected parties, and (ii) $W$, the list of verifiable proofs the party has received (line 1).
  • Upon invocation of the protocol, a party declares two variables: (i) $l$, an index number to access the selected parties one by one from the array $\Sigma$, and (ii) $b$, initially set to zero to indicate that the parties have not reached any agreement (lines 2-3).
  • While loop: Parties are chosen one after another according to the permutation $\Sigma$ of $\{1,\ldots,f+1\}$. Let $l$ denote the index of the party selected in the current loop (the selected party $p_{l}$ is called the candidate). Each party $p_{i}$ follows these steps for the candidate $p_{l}$ (lines 4-22):
    • Broadcasts a $vote$ message to all parties containing $u_{l}=1$ if party $p_{i}$ has received $p_{l}$’s proposal and verifiable proof (including the proposal in the vote), and $u_{l}=0$ otherwise.
    • Waits for $\langle n-f\rangle$ $vote$ messages but does not count votes indicating $u_{l}=1$ unless a valid proposal from $p_{l}$ has been received—either directly or included in the $vote$ message.
    • Runs a binary validated Byzantine agreement biased towards $1$ (see Algorithm 6) to determine whether $p_{l}$ has properly broadcast a valid proposal. Vote $1$ if $p_{i}$ has received a valid proposal from $p_{l}$ and adds the protocol message that completes the verifiable broadcast of $p_{l}$’s proposal to validate this vote. Otherwise, if $p_{i}$ has received $n-f$ $vote$ messages containing $u_{l}=0$, then vote 00; no additional information is needed. If the agreement loop decides $1$, exit the loop.
  • Upon reaching an agreement, a party returns the threshold signature of the agreed proposal.
  • Upon receiving $\langle ID,vote,party,u_{l},m\rangle$, a party checks whether $u_{l}=1$. If it receives $u_{l}=1$, then it assigns $m_{l}=m$.
1
[…]
3      $\Sigma\leftarrow\{\}$
4      $msg\leftarrow createMsg\langle ID,requests\rangle$
5      multi-cast $\langle msg\rangle$
6      wait until $|\Sigma|=n-f$
7      return $\sigma\leftarrow CombineShare_{id}\langle requests,\Sigma\rangle$
8
9
10
[…]
12           if $VerifyShare_{k}\langle m,(k,\sigma_{k})\rangle$  then
13                $\Sigma\leftarrow{\sigma_{k}\cup\Sigma}$
14
15
16                upon receiving $\langle msg\rangle$ from the party $p_{j}$ for the first time do
17                     if $p_{j}\in PParties$  then
18                          $requests\leftarrow msg.requests$
19                          $\sigma_{id}\leftarrow SigShare_{id}\langle sk_{id},requests\rangle$
20                          $reply\langle\sigma_{id}\rangle$
21
22
23
24
25
26                          Local variables initialization:
27                               $\Sigma\leftarrow\{\}$
28
29
30                               upon $Permutation\langle id,instance\rangle$ invocation do
31                                    $\sigma_{i}$leftarrow $CShare\langle r_{id}\rangle$
32                                    $multi-cast\langle SHARE,id,\sigma_{id},instance\rangle$
33                                    wait until $|\Sigma|=2f+1$
34
35                                   return $CToss\langle r_{id},\Sigma\rangle$
36
37
38                                    upon receiving $\langle SHARE,id,\sigma_{k},instance\rangle$ from a party $p_{k}$ for the first time do
39                                         if $CShareVerify\langle r_{k},k,\sigma_{k}\rangle=true$  then
40                                              $\Sigma\leftarrow{\sigma_{k}\cup\Sigma}$
41
42
43
44
Algorithm 4 Protocol for party $p_{i}$
1
[…]
3      $l\leftarrow 0$
4      $b\leftarrow 0$
5      while b=0 do
6           $\Sigma_{vote}\leftarrow\{\}$
7           $l\leftarrow l+1$
8           $party\leftarrow\Sigma\langle l\rangle$
9           $u_{l}\leftarrow 0$
10           $proposal_{l}\leftarrow\perp;$ where $11\leq l\leq f+$ 1
11
12          if $l\in W$ then
13
14               $u_{l}\leftarrow 1$
15                $m_{l}\leftarrow W[l]$
16                multi-cast $\langle ID,vote,party,u_{l}=1,m_{l}\rangle$
17
18                else
19                     multi-cast $\langle ID,vote,party,u_{l}=0,\perp\rangle$
20
21
22                    wait until $|\Sigma_{vote}|$ $=n-f$
23
24
25                    if $u_{l}=1$ then
26                          $v\leftarrow\langle 1,m_{l}\rangle$
27
28                          else
29                               $v\leftarrow\langle 0,\perp\rangle$
30
31                               $\langle b,\sigma\rangle\leftarrow ABBA\langle v\rangle$ biased towards $1$
32
33                               return $\langle b,\sigma\rangle$
34
35
36                              upon receiving $\langle ID,vote,party,u_{l},m\rangle$ for the first time from party $p_{k}$ do
37                                    if $u_{l}=1$  then
38                                         $u_{l}\leftarrow 1$
39                                         $m_{l}\leftarrow m$
40
41                                        $\Sigma_{vote}\leftarrow\Sigma_{vote}+1$
42
43
Algorithm 5 Protocol for party $p_{i}$
Appendix D Agreement protocol
D.1 Asynchronous Binary Byzantine Agreement (ABBA)
The ABBA protocol allows parties to agree on a single bit $b\in\{0,1\}$ [27, 36, 2]. We have adopted the ABBA protocol from [3], as given in Algorithm 6. The expected running time of the protocol is $O(1)$, and it completes within $O(k)$ rounds with probability $1-2^{-k}$. Since the protocol uses a common coin, the total communication complexity becomes $O(kn^{2})$. For more information on how to realize a common coin from a threshold signature scheme, we refer interested readers to the paper [5].
1 upon receiving $b_{input}$, set $est_{0}:=b_{input}$ and proceed as follows in consecutive epochs, with increasing rounds $r$ : do
2
3     multi-cast $BVAL_{r}(est_{r})$
4      bin_valuesr := {}
5      upon receiving $BVAL_{r}(b)$ messages from f+1 parties do
6           if $BVAL_{r}(b)$ has not been sent then
7                multi-cast $BVAL_{r}(b)$
8
9
10
11               upon receiving $BVAL_{r}(b)$ messages from 2f+1 parties do
[…]
13
14
15                    while $bin\_values_{r}\neq\phi$ do
16
17                         multi-cast $AUX_{r}(v)$, where $v\in bin\_values_{r}$
18                          wait until at least (n-f) $AUX_{r}$ messages have been received, such that the set of values carried by these messages,$vals$ are a subset of $bin\_values_{r}$.
19                          $S\leftarrow Coin_{r}.GetCoin()$ see [5, 16]
20                          if vals = {b} then
21                               $est_{r+1}:=b$
22
23                               if (b = S%2) then
24                                    output b
25
26
27                                    else
28                                         $est_{r+1}:=S\%2$
29
30
31                                         continue looping until both a value b is output in some round r, and the value $Coin_{r^{\prime}}=b$ for some round $r^{\prime}>r$.
32
Algorithm 6 ABBA: protocol for party $p_{i}$

### v2

@@
 Abstract
-The multi-valued byzantine agreement protocol (MVBA) in the authenticated setting has been widely used as a core to design atomic broadcast and fault-tolerant state machine replication protocols in asynchronous networks.
-Originating from the seminal work of Cachin et al.
-[13], subsequent research endeavors have sought to optimize protocol efficiency in terms of communication complexity.
-Notable advancements following Cachin’s contributions include: i) VABA [23], requiring multiple protocol instances to achieve agreement on a party’s request, and ii) Dumbo-MVBA [29], employing a cryptographic asynchronous dispersal and recovery methods to manage communication complexity alongside additional computational and communication rounds overheads.
-Our objective is to devise an MVBA protocol that achieves agreement in each instance without extra computation and communication rounds while maintaining the optimal metrics.
-Through meticulous research, we analyze the network/adversary’s message distribution patterns and observe how a party’s message reception varies with the number of request senders.
-We ascertain that opting for a subset of parties ($f+1$, ensuring the inclusion of at least one honest party, at most $f$ number of parties can be faulty ) rather than the entire set ($n$) can facilitate the design of a protocol meeting the aforementioned criteria.
-Central to our design approach is the introduction of the committee in the classic MVBA protocol, wherein a randomly selected subset of ($f+1$, where $n=3f+1$) parties get selected and simultaneously broadcast their requests (transactions) to gather verifiable proofs.
-Successive distributions of these proofs afford us the necessary properties to employ the asynchronous binary Byzantine agreement (ABBA) protocol for reaching an agreement on a selected party’s requests.
-By integrating the committee and ABBA protocols, we devise the optimal MVBA protocol, termed pMVBA (Prioritized-MVBA).
-This protocol exhibits resilience to tolerate up to $\lfloor\frac{n}{3}\rfloor$ Byzantine failures, with an expected runtime of $O(1)$, optimal message complexity of $O(n^{2})$, and optimal communication complexity $O((l+\lambda)n^{2})$ (where $n$ represents the number of parties, $l$ denotes the bit length of the input, and $\lambda$ signifies the security parameter).
+Multi-valued Byzantine agreement (MVBA) protocols are critical components in designing atomic broadcast and fault-tolerant state machine replication protocols in asynchronous networks.
+While these protocols have seen significant advancements, challenges remain in optimizing their communication and computation efficiency without sacrificing performance.
+In this paper, we address the challenge of achieving agreement in MVBA without incurring extra computation and communication rounds.
+Our approach leverages an analysis of message distribution patterns in asynchronous networks, observing that a subset of $f+1$ parties (where at least one party is honest), can achieve an agreement more efficiently than relying on all $n$ parties, where $n=3f+1$, $f$ is the maximum number of faulty parties.
+We introduce a novel protocol, Prioritized-MVBA (pMVBA), which integrates a committee-based selection process and the asynchronous binary Byzantine agreement (ABBA) protocol.
+In this design, a randomly selected subset of $f+1$ parties broadcast their requests, collect verifiable proofs, and utilize these proofs within the ABBA framework to reach an agreement.
+The proposed pMVBA protocol is resilient to up to $\lfloor\frac{n}{3}\rfloor$ Byzantine failures and achieves optimal performance, with an expected runtime of $O(1)$, message complexity of $O(n^{2})$, and communication complexity of $O((l+\lambda)n^{2})$ (similar to state-of-the-art protocols), where $n$ is the number of parties, $l$ is the input bit length, and $\lambda$ is the security parameter.
+The protocol attains $68\%$ higher throughput and $18\%$ lower latency than the classic MVBA protocol (the most efficient protocol).
 Keywords Distributed Systems $\cdot$ Blockchain $\cdot$ Consensus Protocols
@@
 These protocols ensure agreement among parties, even in the presence of malicious actors, making them crucial for maintaining the integrity and functionality of distributed systems.
-The renewed interest in BA protocols has been driven by the success of Bitcoin [35] and other decentralized applications [44], which rely heavily on these protocols for achieving agreement [14].
-Despite the existence of practical BA protocols [1, 33, 5, 37, 41, 31], their reliance on time parameters for ensuring liveness poses challenges in designing agreement protocols.
-The FLP impossibility result [32] demonstrates that deterministic BA protocols cannot guarantee agreement in asynchronous networks.
-This necessitates the development of efficient asynchronous Byzantine Agreement protocols [23, 42, 11, 5, 29, 10, 39, 45, 22] that can overcome the limitations of existing approaches, particularly in terms of communication complexity.
+The renewed interest in BA protocols has been driven by the success of Bitcoin [SATOSHI08] and other decentralized applications [SC], which rely heavily on these protocols for achieving agreement [CACHIN17].
+Despite the existence of practical BA protocols [ABRAHAM20, BYZ08, HONEYBADGER01, PASS17, SHRESTHA20, Hotstuff01], their reliance on time parameters for ensuring liveness poses challenges in designing agreement protocols.
+The FLP impossibility result [CONS03] demonstrates that deterministic BA protocols cannot guarantee agreement in asynchronous networks.
+This necessitates the development of efficient asynchronous Byzantine Agreement protocols [BYZ17, DUAN18, FASTERDUMBO, HONEYBADGER01, LU20, SPEEDINGDUMBO, Joleton, BOLTDUMBO, DAG] that can overcome the limitations of existing approaches, particularly in terms of communication complexity.
 Achieving agreement in asynchronous networks efficiently remains a significant challenge.
-Classic Multi-Valued Byzantine Agreement (MVBA) protocol like Cachin’s [13], while effective, suffer from high communication complexity, often involving terms like $O(n^{3})$.
+Classic Multi-Valued Byzantine Agreement (MVBA) protocol like Cachin’s [CACHIN01], while effective, suffer from high communication complexity, often involving terms like $O(n^{3})$.
 This high complexity is impractical for large-scale systems, necessitating new approaches to reduce communication overhead while ensuring optimal resilience and correctness.
 Cachin et al.
-[13] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision.
+[CACHIN01] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision.
 Cachin’s MVBA protocol provides polynomial-time complexity but suffers from high communication overhead, with a complexity of $O(n^{3})$.
-VABA [23] introduced a view-based approach to eliminate the $O(n^{3})$ term, while Dumbo-MVBA [29] used erasure codes to handle large input sizes more efficiently, reducing the communication complexity to $O(ln+\lambda n^{2})$.
+VABA [BYZ17] introduced a view-based approach to eliminate the $O(n^{3})$ term, while Dumbo-MVBA [LU20] used erasure codes to handle large input sizes more efficiently, reducing the communication complexity to $O(ln+\lambda n^{2})$.
 However, these optimizations often introduced additional rounds of communication and computation.
@@
 These enhancements make the pMVBA protocol suitable for large-scale decentralized applications, addressing the scalability challenges faced by traditional MVBA protocols.
-The main contributions of this paper are as follows:
-  • We introduce a novel committee selection protocol that ensures the presence of at least one honest party in the selected subset, enhancing the protocol’s resilience against Byzantine failures.
-  • We propose the pMVBA protocol, which reduces the communication complexity from $O(n^{3})$ to $O((l+\lambda)n^{2})$ by dynamically selecting a subset of parties to broadcast their proposals.
-  • Our protocol integrates the ABBA protocol to achieve agreement with probability $1$ on the proposals broadcast by committee members, ensuring robust agreement mechanisms.
-  • We provide a thorough theoretical analysis demonstrating that the pMVBA protocol achieves an expected constant asynchronous round count while maintaining optimal resilience and correctness.
-  • We present extensive case studies and performance evaluations, showing that the pMVBA protocol is effective and scalable for large-scale decentralized applications, outperforming traditional MVBA protocols in terms of communication overhead and efficiency.
+(a) O(n) bottleneck
+(b) No bottleneck
+Figure 1: Comparison of two protocols.
+We show the time complexity of each sub-protocol and the step where our protocol wins in terms of communication.
+The txtx represents the transaction, the tsts represents the threshold signature.
+$1,1,1,0$ represents the bit array and each color for the particular party and its messages.
+The main contributions of this paper are as follows: The main contributions of this project are as follows:
+  • We propose the pMVBA protocol, which reduces the communication complexity from $O(n^{3})$ [SECURE03] to $O((l+\lambda)n^{2})$ by dynamically selecting a subset of parties using the cryptographic coin-tossing technique to broadcast their proposals.
+  • We introduce a committee in the classical MVBA protocol [SECURE03] that ensures that the protocol reaches an agreement with probability $1$, unlike other committee-based protocols.
+We replace the $O(n)$ bit Commit step of the classic protocol with $O(1)$ Recommend step.
+  • We integrate the ABBA biased towards $1$ protocol to mitigate the effect of the reduction in the total number of proposals due to the introduction of the committee.
+  • We present proof, extensive case studies, and experimental results to show that our approach (committee and Recommend step) fills the requirement of the ABBA protocol to reach an agreement on a committee member’s proposal.
+  • We implement pMVBA and extensively test it among $n=4$, 77 or $10$ nodes in gRPC based simulator, making detailed Comparison to the state-of-the-art MVBA protocols, including ClassicMVBA[SECURE03] (CKSP01), and one recent result, VABA (Abraham et al.
+[BYZ17]).
+At all system scales, the peak throughput of pMVBA is better than any of the other tested MVBA protocols (up to $180\%$).
+(see Table 1).
+    More importantly, the latency of pMVBA protocol is significantly less than that of others (e.
+g.
+VABA [BYZ17], CKSP01 [SECURE03]).
+Actually, the latency of pMVBA is nearly independent of its throughput, indicating the effectiveness of the protocol to resolve the throughput-latency tension lying in the previous protocols.
+As shown in Figure 2, we see that pMVBA has the lowest increment in terms of latency, and the increment ratio is also lowest among other protocols.
+This result indicates the effectiveness of the proposed MVBA protocol.
+    Table 1: In comparison with existing MVBA protocols.
+    Protocols Peak throughput (tps) Latency at peak-tps(sec)
+    n=4 n=7 n = 10 n=4 n=7 n = 10
+    VABA [BYZ17] 30.5 15.1 9 16.39 33 55.35
+    CKSP01 [SECURE03] 53.47 23.8 12 9.35 21 41.92
+    pMVBA (ours) 85.67 37.87 22 5.84 13.2 22.89
+    Figure 2: Latency increment of MVBA protocols when throughput increases from minimum to maximum (n=10)
 2 Related Work
@@
 The Byzantine Generals Problem, introduced by Lamport et al.
-[26], laid the theoretical foundation for BA protocols, emphasizing the challenges of achieving an agreement in the presence of faulty or malicious nodes.
+[LAMPORT82], laid the theoretical foundation for BA protocols, emphasizing the challenges of achieving an agreement in the presence of faulty or malicious nodes.
 Early BA protocols, such as those by Pease et al.
-[38], focused on synchronous networks where communication occurs in fixed rounds.
+[PEASE80], focused on synchronous networks where communication occurs in fixed rounds.
 However, these protocols often relied on time assumptions for ensuring liveness.
-Practical Byzantine Fault Tolerance (PBFT), proposed by Castro and Liskov [15], was a significant advancement in making BA protocols practical for real-world applications.
+Practical Byzantine Fault Tolerance (PBFT), proposed by Castro and Liskov [CASTRO99], was a significant advancement in making BA protocols practical for real-world applications.
 PBFT introduced an efficient consensus protocol for partially-synchronous networks, ensuring safety and liveness even with up to $f<\frac{n}{3}$ Byzantine faults.
-Several modern BA protocols build upon the principles of PBFT, including HotStuff [46] and Tendermint [12], optimizing various aspects such as communication complexity and responsiveness.
-The FLP impossibility result [18] demonstrated that no deterministic BA protocol could guarantee an agreement in an asynchronous network with a single faulty node.
+Several modern BA protocols build upon the principles of PBFT, including HotStuff [HOTSTUFF19] and Tendermint [TENDERMINT14], optimizing various aspects such as communication complexity and responsiveness.
+The FLP impossibility result [FLP85] demonstrated that no deterministic BA protocol could guarantee an agreement in an asynchronous network with a single faulty node.
 This led to the exploration of probabilistic and randomized approaches for ABA.
-Ben-Or’s protocol [8] was one of the earliest randomized BA protocols, using coin-flipping to break symmetry among nodes.
-Recent protocols like HoneyBadgerBFT [34] and Dumbo [28] have focused on optimizing ABA for high throughput and low latency.
+Ben-Or’s protocol [BENOR83] was one of the earliest randomized BA protocols, using coin-flipping to break symmetry among nodes.
+Recent protocols like HoneyBadgerBFT [MILLER16] and Dumbo [DUMBO20] have focused on optimizing ABA for high throughput and low latency.
 Multi-Valued Byzantine Agreement Protocols.
 Cachin et al.
-[13] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision.
+[CACHIN01] introduced the concept of MVBA, where nodes agree on a value from a large input space rather than a binary decision.
 Cachin’s MVBA protocol provides polynomial-time complexity but suffers from high communication overhead, with a complexity of $O(ln^{2}+\lambda n^{2}+n^{3})$.
-Subsequent works, such as VABA [23] and Dumbo-MVBA [29], aimed to reduce the communication complexity of MVBA.
+Subsequent works, such as VABA [BYZ17] and Dumbo-MVBA [LU20], aimed to reduce the communication complexity of MVBA.
 VABA introduced a view-based approach to eliminate the $O(n^{3})$ term, while Dumbo-MVBA used erasure codes to handle large input sizes more efficiently, reducing the communication complexity to $O(ln+\lambda n^{2})$.
@@
 Committee-based protocols have emerged as a promising approach to reduce communication complexity in BA protocols.
-Algorand [20] introduces the concept of committee members, wherein a predefined set of parties collaborates to reach an agreement.
+Algorand [BYZ21] introduces the concept of committee members, wherein a predefined set of parties collaborates to reach an agreement.
 Algorand operates within a permission-less (non-trusted) setting.
-In a similar vein, COINcidence [19] presents a protocol within a trusted setup, leveraging committee-based participation to achieve agreement with high probability.
+In a similar vein, COINcidence [COINCIDENCE] presents a protocol within a trusted setup, leveraging committee-based participation to achieve agreement with high probability.
 Our approach diverges from both Algorand and COINcidence.
@@
 Furthermore, our protocol guarantees an agreement on a single party’s requests with a probability of 1.
-Conversely, FasterDumbo [11] utilizes the committee concept to minimize the number of components in the agreement process, yielding outputs from $n-f$ parties’ requests.
-Beat [42] amalgamates the most effective elements of HoneyBadgerBFT’s concrete implementation to showcase superior performance across varied settings.
+Conversely, FasterDumbo [FASTERDUMBO] utilizes the committee concept to minimize the number of components in the agreement process, yielding outputs from nfn-f parties’ requests.
+Beat [DUAN18] amalgamates the most effective elements of HoneyBadgerBFT’s concrete implementation to showcase superior performance across varied settings.
 Optimistic Fastlane Mechanisms.
 Certain works, such as those by Victor et al.
-[25] and Cachin et al.
-[17], have explored the incorporation of an optimistic ’fastlane’ mechanism, designed for operation within partially synchronous networks.
+[VICTOR01] and Cachin et al.
+[CACHIN02], have explored the incorporation of an optimistic ’fastlane’ mechanism, designed for operation within partially synchronous networks.
 However, in instances of sluggish agreement processes, these mechanisms transition to asynchronous Byzantine agreement protocols.
-Recent endeavors, such as those by Joleton [39] and Boltdumbo [45], have adopted a similar strategy.
+Recent endeavors, such as those by Joleton [Joleton] and Boltdumbo [BOLTDUMBO], have adopted a similar strategy.
 Our work remains pertinent within this context as we rely on asynchronous protocols when the network exhibits Byzantine behavior, particularly within the pessimistic path.
-Given our protocol’s focus on outputting requests from a single party, we omit comparisons with DAG-rider [22] and Aleph [21], which utilize directed acyclic graphs for agreement alongside sequential ACS.
-Additionally, recent efforts have aimed to eliminate the private setup phase in asynchronous Byzantine agreement protocols, emphasizing asynchronous distributed key generation [40, 30, 7].
+Given our protocol’s focus on outputting requests from a single party, we omit comparisons with DAG-rider [DAG] and Aleph [ALEPH], which utilize directed acyclic graphs for agreement alongside sequential ACS.
+Additionally, recent efforts have aimed to eliminate the private setup phase in asynchronous Byzantine agreement protocols, emphasizing asynchronous distributed key generation [DIS01, DIS02, DIS03].
 3 System Model
-We use the standard notion [16, 13] to describe a distributed algorithm involving $n$ parties and an adversary in an authenticated setting.
+We use the standard notion [SECURE02, CACHIN01] to describe a distributed algorithm involving $n$ parties and an adversary in an authenticated setting.
 This section explains the foundational assumptions, network model, adversary model, computational model, and cryptographic tools used in our protocol.
@@
 The generation and distribution of these keys are out of the scope of this paper.
-We follow standard literature for key generation and distribution and refer interested readers to [9, 4, 6].
+We follow standard literature for key generation and distribution and refer interested readers to [BORN01, BOLD01, Kate01].
 We use the node and party alternatively throughout the paper.
@@
 The adversary can corrupt $f$ parties among the $n$ parties, where $f<\frac{n}{3}$.
-Computational Model: We adopt standard modern cryptographic assumptions and definitions from [16, 13].
+Computational Model: We adopt standard modern cryptographic assumptions and definitions from [SECURE02, CACHIN01].
 The assumptions allow the parties and the adversary to be probabilistic polynomial-time interactive Turing machines.
@@
 This property is crucial for ensuring the protocol’s correctness.
-Therefore, it is called the validated asynchronous byzantine agreement [13, 24, 47].
+Therefore, it is called the validated asynchronous byzantine agreement [CACHIN01, VABA, BYZ20].
 Multi-valued Byzantine Agreement The MVBA protocol provides a polynomial-time computable predicate $Q$.
@@
 Guo et al.
-[11] constructed the $(1,\kappa,\epsilon)$-Committee Selection protocol using a threshold coin-tossing mechanism (see Appendix A.3), which is derived from threshold signatures.
+[FASTERDUMBO] constructed the $(1,\kappa,\epsilon)$-Committee Selection protocol using a threshold coin-tossing mechanism (see Appendix A.3), which is derived from threshold signatures.
 The protocol ensures that at least one honest party is a committee member with overwhelming probability $1-\epsilon-\text{neg}(\lambda)$, where $\text{neg}(\lambda)$ is a negligible function in the cryptographic security parameter $\lambda$, and $\epsilon$ is $exp(-\Omega\kappa)$.
-Cryptographic Abstractions: We design a distributed algorithm in authenticated settings where we use robust, non-interactive threshold signatures to authenticate messages and a threshold coin-tossing protocol to select parties randomly [16].
+Cryptographic Abstractions: We design a distributed algorithm in authenticated settings where we use robust, non-interactive threshold signatures to authenticate messages and a threshold coin-tossing protocol to select parties randomly [SECURE02].
   1. 1.
-    Threshold Signature Scheme: We utilize a threshold signature scheme introduced in [43, 16].
+    Threshold Signature Scheme: We utilize a threshold signature scheme introduced in [THRESH01, SECURE02].
 The basic idea is that there are $n$ parties, up to $f$ of which may be corrupted.
@@
   2. 2.
-    Threshold Coin-Tossing Scheme: The threshold coin-tossing scheme, as introduced in [43, 16], involves parties holding shares of a pseudorandom function $F$ that maps the name $C$ (an arbitrary bit string) of a coin.
+    Threshold Coin-Tossing Scheme: The threshold coin-tossing scheme, as introduced in [THRESH01, SECURE02], involves parties holding shares of a pseudorandom function $F$ that maps the name $C$ (an arbitrary bit string) of a coin.
 We use a distributed pseudorandom function as a coin that produces $k^{\prime\prime}$ random bits simultaneously.
@@
 We describe the security properties of the scheme and related algorithms in Appendix A.3.
-4 The Proposed Protocol
-(a) An overview of the proposed protocol framework, highlighting our contributions: committee selection, the VCBC protocol for committee members, recommendations, and Sequential ABBA.
+4 Design of pMVBA
+4.1 pMVBA Overview
 In this section, we present a protocol for multi-valued Byzantine agreement capable of tolerating up to $f<\frac{n}{3}$ Byzantine faults.
 The protocol features an expected communication bit complexity of $O(ln^{2}+\lambda n^{2})$ and an expected constant asynchronous round count.
-Our modular implementation comprises five distinct sub-protocols: committee selection, prioritized verifiable consistent broadcast protocol (pVCBC), Propose-Recommend, random order/permutation generation, and Sequential-ABBA.
-The framework of the proposed protocol is illustrated in Figure 1(a).
-The violet-colored boxes indicate our contributions to the protocol.
-We break down the protocol into two main parts.
-The top part includes the committee selection and the VCBC protocol for the committee members, while the bottom part handles the distribution of the committee members’ messages and the agreement protocol.
-Section 4.1 provides an overview and the construction of party selection and broadcast (top part of the framework), and Section 4.2 offers a detailed exposition of the protocol, including pseudocode and descriptions.
-4.1 Optimizing Message and Communication Overhead
-To enhance the efficiency of Byzantine agreement protocols, we propose a novel method to minimize the number of broadcast requests.
-Rather than requiring all $n$ parties to broadcast their requests, we select a subset of parties to do so, ensuring that the protocol’s progress and security properties are maintained.
-Party Selection and Broadcast: In traditional asynchronous Byzantine agreement protocols, all $n$ parties are typically allowed to broadcast their proposals.
-However, the main factor contributing to the $O(n^{3})$ communication complexity in classic MVBA protocols is that each of the $n$ parties must broadcast an $O(n)$-size array.
-These arrays include broadcast information from all parties.
-Our approach aims to reduce the number of broadcasting parties, thereby eliminating the need for transmitting an $O(n)$-size array and decreasing the number of required Asynchronous Binary Byzantine Agreement (ABBA) instances.
-From this point forward, we use proposal instead of request and propose instead of broadcasting requests.
-To achieve this reduction, we select at least $f+1$ parties to broadcast their proposals, ensuring that there is at least one honest party among them.
-This selection is crucial for guaranteeing the protocol’s progress.
-We employ a committee selection protocol similar to FasterDumbo [11], which uses a cryptographic coin-tossing scheme to randomly select $\kappa$ (a security parameter) parties.
-We set $\kappa=f+1$ to ensure the inclusion of at least one honest party, increasing the likelihood of multiple honest parties being selected.
-Dynamic selection of parties for each protocol instance prevents adversaries from corrupting fixed parties, ensures fair opportunity for all parties to propose, and mitigates the risk of starvation and Denial-of-Service (DoS) attacks.
-The committee selection process must be immune to adversary influence.
-The cryptographic coin-tossing scheme ensures random selection, requiring honest party participation after the completion of the previous instance.
-We used the committee member/selected party alternatively.
+Our modular implementation comprises five distinct sub-protocols: Committee Selection (CS), Prioritized Verifiable Consistent Broadcast (pVCBC), Propose-Recommend, Random Order/Permutation Generation, and Sequential-ABBA.
+The framework of the proposed protocol is illustrated in Figure LABEL:fig:Framework.
+4.2 Committee Selection
+In the proposed pMVBA protocol, the CS sub-protocol plays a critical role in optimizing efficiency by reducing the communication complexity inherent in Byzantine agreement protocols.
+Rather than requiring all $n$ parties to broadcast their proposals, our method strategically selects a subset of at least $f+1$ parties, based on the party ID and the current instance of the protocol, to perform this task.
+This selection ensures the protocol’s progress while maintaining its security properties, with the output being a set of $f+1$ parties designated as committee members for that specific instance.
+To achieve this, the CS protocol utilizes a cryptographic coin-tossing scheme, inspired by Dumbo [FASTERDUMBO], to dynamically and randomly select $\kappa=f+1$ parties for each protocol instance.
+This selection process guarantees that at least one honest party is included in the committee, thereby preserving the integrity of the protocol.
+The dynamic nature of party selection mitigates several risks, including adversarial corruption, ensuring fair participation across all parties and reducing the likelihood of starvation and Denial-of-Service (DoS) attacks.
+The randomness introduced by the cryptographic coin-tossing scheme is crucial in preventing adversary influence and securing the CS process.
+The CS protocol is illustrated in Algorithm 1 and involves the following steps:
+  • Generating Coin-Shares: Upon invocation, each party generates a coin-share $\sigma_{i}$ and broadcasts it to all parties. The party then waits to receive $f+1$ coin-shares. (lines 3-5)
+  • Verifying Coin-Shares: Upon receiving a coin-share from another party $p_{k}$, the receiving party verifies the coin-share and adds it to the set $\Sigma$ until $f+1$ valid shares are collected.
+(lines 8-10)
+  • Selecting Parties: After collecting $f+1$ valid coin-shares, the $CToss$ function is used to determine the selected subset of parties. (lines 6-7)
+1
+2Local variables initialization:
+3   $\Sigma\leftarrow\{\}$
+4   $CUR\_INSTANCE\leftarrow 0$
+5
+6
+7   upon $CS\langle ID\rangle$ invocation do
+8     $CUR\_INSTANCE\leftarrow ID.INSTANCE$
+9     $id\leftarrow ID.id$
+10     $\sigma_{id}$ $\leftarrow$ $CShare\langle id\rangle$
+11
+12    send $\langle SHARE,id,\sigma_{id},CUR\_lINSTANCE\rangle$ to all parties
+13     wait until $|\Sigma|=f+1$
+14     return $CToss\langle id,\Sigma\rangle$
+15
+16
+17    upon receiving $\langle SHARE,id,\sigma_{k},INSTANCE\rangle$ from a party $p_{k}$ for the first time do
+18       if CUR_INSTANCE=INSTANCECUR\_INSTANCE=INSTANCE and $CShareVerify\langle id,k,\sigma_{k}\rangle=true$ then
+19         $\Sigma\leftarrow\Sigma\cup\{\sigma_{k}\}$
+20
+21
+Algorithm 1 Committee Selection for party $p_{i}$
 (a) Committee-Selection protocol.
-Each party shares their $s_{i}=CShare(r_{i})$ and a party requires f+1 shares.
-We see a party use $CToss(r_{1},s)$ function to get the selected parties and each party gets the same set of parties.
-We show only for the two parties but it is true for every party.
+Each party shares their $s_{i}=CShare(r_{i})$, and a party requires $f+1$ shares.
+The $CToss(r_{1},s)$ function is used to select the parties, ensuring that each party gets the same set.
+This illustration shows the process for two parties, but it applies universally.
 (b) pVCBC protocol.
-Each selected party (here $p_{1}$ and $p_{2}$ ) propose its requests.
-Each receiving party verify the requests and add its sign-share on the message and returns to the sender.
+Each selected party (here $p_{1}$ and $p_{2}$ ) proposes its requests.
+Each receiving party verifies the requests, adds its sign-share to the message, and returns it to the sender.
 A party waits for 2f+1 sign-shares (here it is 3 because f = 1).
 The 2f+1 sign shares are necessary to create a threshold-signature ($ts1$).
-Committee Selection Protocol: The committee selection protocol (CS) is an integral part of our approach.
-It ensures that a reliable subset of parties is chosen to broadcast proposals, thus reducing communication complexity.
-The steps of the CS protocol are illustrated in Algorithm 1 and described as follows:
-  • Generating Coin-Shares: Upon invocation, a party generates a coin-share $\sigma_{i}$ for the current instance and broadcasts it to all parties, waiting to receive $f+1$ coin-shares. (lines 3-5)
-  • Verifying Coin-Shares: When a party receives a coin-share for the first time from a party $p_{k}$, it verifies the coin-share and accumulates it in set $\Sigma$.
-This process continues until $f+1$ valid shares are collected.
-(lines 8-10)
-  • Selecting Parties: After receiving $f+1$ valid coin-shares, a party uses the $CToss$ function along with the collected coin-shares to select $f+1$ parties. (lines 6-7)
+Figure 3: Committee Selection and pVCBC protocol
+The CS protocol, as depicted in Figure 3(a), ensures a secure and efficient selection of broadcasting parties, which significantly reduces the overall communication complexity of the pMVBA protocol.
+Definitions of the cryptographic coin-tossing functions (e.g., CShareVerifyCShareVerify) are provided in Appendix A.3.
+4.3 pVCBC
+Following the selection of committee members through the Committee Selection protocol, each selected member must provide a verifiable proof of their proposal to ensure that it has been broadcast to at least $f+1$ honest parties.
+Specifically, the input for this protocol includes the ID, requests, and the selected parties, while the output is a threshold-signature—a verifiable proof that the same requests have been sent to at least $f+1$ honest parties.
+This proof is crucial for maintaining the integrity and consistency of the protocol, as it verifies that the proposal has been correctly disseminated among the parties.
+Traditionally, the Verifiable Consistent Broadcast (VCBC) protocol is used to generate such proofs, enabling every party to produce a verifiable record of their broadcast proposals.
+The definition of verifiability and the VCBC protocol itself are detailed in Appendix A.1.
+However, since our protocol restricts broadcasting to only the selected committee members, we utilize a slightly modified version of the VCBC protocol, which we term pVCBC (Prioritized Verifiable Consistent Broadcast).
+The pVCBC protocol aligns with the selective broadcasting approach from the Committee Selection process.
+This adaptation ensures that when a party receives verifiable proof from another, there’s no need to verify its origin from a selected committee member, as the protocol inherently guarantees this.
+This streamlines verification and reduces unnecessary checks, preserving the efficiency introduced by Committee Selection.
+The pVCBC protocol’s construction is detailed in Algorithm 2, and an illustration is provided in Figure 3(b), visually representing its interactions and verification steps as a critical component of the overall pMVBA protocol.
+Construction of the pVCBC
+  • Upon invocation of the pVCBC protocol, a party sends (ID, requests, SelectedParties) to the parties. (lines 03).
+  • Upon receiving a message (ID, requests, SelectedParties) from a party $p_{k}$, a party checks whether the sender is a selected party.
+If the sender is a selected party, the party adds its signature share (signsharesign-share) to the message, resulting in $\sigma$.
+The party then replies with $\sigma$ to the sender (lines 11-14).
+  • Upon receiving a signature share $\sigma_{k}$ from a party $p_{k}$, a selected party adds the signature share $\sigma_{k}$ to its set $\Sigma$ (lines 07 - 09).
+  • A selected party waits for $\langle n-f\rangle$ valid signature shares.
+Upon receiving nfn-f signature shares, the party combines them to generate a threshold signature $\rho$ (proof that the party has sent the same request to at least $f+1$ honest parties) and returns $\rho$ to the caller (line 05).
 1
+[…]
+3   $\Sigma\leftarrow\{\}$
+4   send $\langle ID,requests\rangle$ to all parties
+5   wait until $|\Sigma|=n-f$
+6   return $\rho\leftarrow CombineShare_{id}\langle requests,\Sigma\rangle$
+7
+8
+9
+[…]
+11     if $VerifyShare_{k}\langle requests,(k,\sigma_{k})\rangle$ then
+12       $\Sigma\leftarrow{\Sigma\cup\{\sigma_{k}\}}$
+13
+14
+15       upon receiving $\langle ID,requests\rangle$ from the party $p_{j}$ for the first time do
+16         if $p_{j}\in SelectedParties$ then
+17           $\sigma_{id}\leftarrow SigShare_{id}\langle sk_{id},requests\rangle$
+18           $reply\langle requests,\sigma_{id}\rangle$
+19
+20
+21
+22
+23
 24           Local variables initialization:
 25             $\Sigma\leftarrow\{\}$
-4
-5
-6      upon $CS\langle id,instance\rangle$ invocation do
-7           $\sigma_{i}$leftarrow $CShare\langle r_{i}\rangle$
-8           $multi-cast\langle SHARE,id,\sigma_{id},instance\rangle$
-9           wait until $|\Sigma|=f+1$
-10           return $CToss\langle r_{i},\Sigma\rangle$
-11
-12
+26
+27
+28             upon $RandomOrder\langle ID,SelectedParties\rangle$ invocation do
+29               $\sigma_{i}$ $\leftarrow$ $CShare\langle r_{id}\rangle$
+30               $id=ID.id$
+31               $\textbf{send}\langle SHARE,id,\sigma_{id},ID.INSTANCE\rangle$ to all parties
+32               wait until $|\Sigma|=2f+1$
+33
+34              return filterSelectedParties( $CToss\langle r_{id},\Sigma\rangle,SelectedParties$)
+35
+36
 37               upon receiving $\langle SHARE,id,\sigma_{k},INSTANCE\rangle$ from a party $p_{k}$ for the first time do
 38                 if $CShareVerify\langle r_{k},k,\sigma_{k}\rangle=true$ then
-15                     $\Sigma\leftarrow{\sigma_{k}\cup\Sigma}$
-16
+39                   $\Sigma\leftarrow{\Sigma\cup\{\sigma_{k}\}}$
+40
+41
+42
+Algorithm 2 pVCBC and RandomOrder protocols for party $p_{i}$
+4.4 Propose-Recommend
+In the pMVBA protocol, the Propose-Recommend step replaces the traditional commit step found in the classical MVBA protocol.
+This shift is central to enhancing the protocol’s efficiency by significantly reducing complexity.
+While the classic commit step involves collecting nfn-f verifiable proofs and creating an array of length $n$, our recommend step simplifies the process by broadcasting verifiable proofs directly.
+The propose step remains largely similar to the one in the classic protocol, but with a key difference: only the selected parties, as determined by the CS process, are responsible for proposing their requests.
+The input for this step includes a proposal and a threshold-signature, but it does not produce any direct output.
+The recommend step then follows, where the input consists of the proposal and the threshold-signature.
+The output of this step is a list of threshold-signatures (collected from the nfn-f recommendation message), ensuring that at least one verifiable proof reaches the majority of parties.
+This process is critical for moving the protocol towards an agreement.
+The recommend step involves the following key actions:
+  • Initialization: A party initializes an empty set $\Sigma$. (line 2)
+  • Broadcast Recommendation: A party creates a recommendation message and broadcasts it to all parties. (line 6)
+  • Recommendation Collection: Upon receiving a recommendation, a party verifies the threshold-signature and instance.
+If valid, the party adds the proposal and $\sigma$ to its set $\Sigma$.
+(lines 11-14)
+  • A party waits for nfn-f recommendation messages before proceeding. (line 6-7)
+The detailed SCR (Send, Collect, Recommend) protocol is shown in Algorithm 3, which formalizes the steps involved in the recommend process.
+1Local variables initialization:
+2 $\Sigma\leftarrow\{\}$
+3 $CUR\_INSTANCE\leftarrow 1$
+4 $rCount\leftarrow 0$
+5
+6 upon $SCR\langle ID,ID_{k},Proposal,\rho\rangle$ invocation do
+7 send $\langle RECOMMENDATION,ID_{k},Proposal,\rho\rangle$ to all parties
+8 $CUR\_INSTANCE\leftarrow ID.INSTANCE$
+9 $rCount\leftarrow 0$
+10 wait until rCount=nfrCount=n-f
+11 return $\Sigma$
+12
+13
+14
+15upon receiving $\langle RECOMMENDATION,ID_{j},Proposal,\rho\rangle$ from party $p_{j}$ for the first time do
+16 if ID.INSTANCE=CUR_INSTANCEID.INSTANCE=CUR\_INSTANCE $\wedge$ $Verify_{id}\langle Proposal,\rho\rangle=true$ $\wedge ID_{j}.id\notin\Sigma$ then
+17    $\Sigma\leftarrow{\Sigma\cup\{ID_{j}.id:\langle Proposal,\rho\rangle\}}$
+18
+19    $rCount\leftarrow rCount+1$
+20
+21
+22
+23
+Algorithm 3 Recommend Protocol for all parties
+4.5 Random Order
+Following the Propose-Recommend step, the next crucial phase in the pMVBA protocol is to determine the order in which the selected parties will proceed.
+The purpose of this Random Order step is to generate a random permutation of the selected parties, ensuring that the process remains fair and secure.
+This step is particularly important for preventing adversaries from manipulating the message delivery order, which could potentially increase the number of asynchronous rounds required to reach an agreement.
+The input for this step includes the party ID and the current protocol instance, and the output is a permutation of the selected parties.
+The random order is generated after the distribution and receipt of the threshold number of recommendation messages, ensuring the process is not vulnerable to adversarial interference.
+The Random Order protocol involves the following steps :
+  • Coin-Share Generation: Upon invocation of the RandomOrder protocol, a party generates a coin-share $\sigma_{i}$ for the current instance and broadcasts it to every party.
+The party then waits for $2f+1$ coin-shares.
+(lines 19-24)
+  • Coin-Share Verification: When a party receives a coin-share from another party $p_{k}$ for the first time, it verifies the coin-share, ensuring it is from $p_{k}$, and accumulates the coin-share in the set $\Sigma$.
+The party continues to respond to coin-shares until it has received $2f+1$ valid shares.
+(lines 27-28)
+  • Permutation Generation: Once $2f+1$ valid coin-shares have been received, the party uses its $CToss$ function and the collected coin-shares to generate a permutation of the $n$ parties and filtered the selected parties and return.
+(line 24 of Algorithm 2).
+This process is formalized in the pseudocode for the random-order generation protocol, as shown in Algorithm 2.
+By generating a random order of the selected parties, this step ensures that the protocol remains robust against adversarial manipulation and maintains the efficiency of the agreement process.
+4.6 Sequential-ABBA
+The final phase in the pMVBA protocol is the Sequential-ABBA protocol, which is responsible for reaching an agreement on one of the proposals submitted by the selected parties.
+Building on the random order generated in the previous step, the Sequential-ABBA protocol runs an agreement loop that systematically evaluates each selected party’s proposal.
+The input to this protocol consists of the permutation list generated in the Random Order step and the list of recommendations obtained from the Propose-Recommend step.
+The output is the final proposal agreed upon by the parties.
+In Sequential-ABBA, we employ an asynchronous binary Byzantine agreement protocol that is biased towards $1$, allowing for efficient agreement on proposals.
+The protocol iterates through the permutation list, running the agreement process for each selected party until an agreement is achieved on a single proposal.
+This approach ensures that the final decision is reached in a manner that is both robust and fair, in line with the protocol’s overall design.
+The detailed construction of the Sequential-ABBA protocol is provided below, where the specific steps and mechanisms are described.
+This protocol serves as the culminating step in the pMVBA process, ensuring that all parties converge on a single, agreed-upon proposal, thus completing the agreement procedure.
+Construction of the Sequential-ABBA
+Since the Sequential-ABBA protocol guarantees that the parties will eventually reach an agreement on one of the selected party’s proposals, the protocol runs a loop for the selected parties until it agrees on a valid proposal.
+The pseudocode of the Sequential-ABBA protocol is given in Algorithm 4, and a step-by-step description is provided below:
+  • The Sequential-ABBA protocol takes two arguments: (i) PermutationList, the permutation of the selected parties, and (ii) RecommendationList, the list of verifiable proofs and proposals the party has received (line 1).
+  • Upon invocation of the protocol, a party declares two variables: (i) indexindex, an index number to access the selected parties one by one from the array PermutationList, and (ii) bitbit, initially set to zero to indicate that the parties have not reached any agreement (lines 2-3).
+  • While loop: Parties are chosen one after another according to the permutation PermutationList of $\{1,\ldots,f+1\}$.
+Let indexindex denote the index of the party selected in the current loop (the selected party $p_{index}$ is called the candidate).
+Each party $p_{i}$ follows these steps for the candidate $p_{index}$ (lines 4-25):
+    • Broadcasts a $VOTE$ message to all parties containing $u_{index}=1$ if party $p_{i}$ has received $p_{index}$’s proposal and verifiable proof (including the proposal in the VOTE or from RecommendationList), and $u_{index}=0$ otherwise.
+(lines 12-18)
+    • Waits for $\langle n-f\rangle$ VOTEVOTE messages (line 19) but does not count votes indicating $u_{index}=1$ unless a valid proposal from the party $p_{id}$ has been received—either directly or included in the $VOTE$ message (lines 21 - 22).
+    • Runs a binary asynchronous Byzantine agreement biased towards $1$ (see Algorithm 6) to determine whether $p_{index}$ has properly broadcast a valid proposal.
+Vote $1$ if $p_{i}$ has received a valid proposal from $p_{index}$ through RecommendationList and add the protocol message that completes the verifiable broadcast of $p_{index}$’s proposal to validate this vote.
+Otherwise, if $p_{i}$ has received nfn-f VOTEVOTE messages containing $u_{index}=0$, then vote 0; no additional information is needed.
+If the agreement loop decides $1$, exit the loop.
+(lines 21-25)
+  • Upon reaching an agreement, if the proposal is empty, then use the threshold signature $\rho$ to get the proposal from the selected party and return the proposal. (lines 26-28)
+  • Upon receiving $\langle ID,VOTE,party,u_{index},m\rangle$, a party checks whether $u_{index^{\prime}}=1$. If it receives $u_{index^{\prime}}=1$, then it assigns $m_{index}=m$. (lines 29-33)
+1
+[…]
+3   $index\leftarrow 0$
+4   $bit\leftarrow 0$
+5
+6  while bit=0 do
+7     $\Sigma\leftarrow\{\}$
+8
+9    $index\leftarrow index+1$
+10
+11    $proposal\leftarrow\perp$
+12     $\rho\leftarrow\perp$
+13
+14    $id\leftarrow PermutationList[index]$
+15     $u_{index}\leftarrow 0$
+16     $m_{index}\leftarrow<proposal,\rho>$
 17
-Algorithm 1 Committee Selection Protocol for party $p_{i}$
-We provide a illustration of the CS protocol in figure 2(a) and the cryptographic coin-tossing definitions (e.g. $CShareVerify$) in the Appendix A.3.
-Broadcast protocol A committee member requires a verifiable proof of its proposal to demonstrate that it has broadcast the same proposal to at least $f+1$ honest parties.
-Typically, the VCBC is employed for this purpose, allowing every party to generate a verifiable proof.
-Since our protocol restricts proposal broadcasting to only the selected parties, we employ a slightly modified version of the verifiable consistent broadcast protocol, termed pVCBC.
-This adaptation ensures that if a party receives a verifiable proof from another party, it does not need to verify whether the proof originates from a selected party or not.
-The definition of the VCBC protocol and the construction of the pVCBC protocol can be found in the Appendix A.1 and Appendix C.
-We have also provided an illustration of the protocol in Figure 2(b).
-4.2 Prioritized MVBA
-This section details the MVBA module in an asynchronous network.
-We propose a novel MVBA protocol called Prioritized-MVBA (pMVBA), which uses a committee as a prioritized parties and reduces the communication complexity of the classic MVBA protocol.
-Figure 3: New-VBAConst protocol.
-After completing the pVCBC broadcast, each party proposes its value along with verifiable proof.
-Upon receiving verifiable proof, the party recommends the received proposal.
-Each party waits for $n-f$ recommendations, then uses the coin-tossing scheme (CShare) to generate a permutation and initiates the ABBA protocol to reach an agreement on one of the proposals.
-Overview of pMVBA The pMVBA protocol is inspired by the asynchronous binary Byzantine agreement methodology from Cachin et al.
-[13].
-It significantly reduces the number of initial broadcasts, eliminating costly sub-components by leveraging the message distribution pattern of the adversary.
-In classic MVBA, each party broadcasts an $O(n)$-size array to achieve $O(1)$ expected constant runtime.
-Our pMVBA protocol as shown in Fig 3 and explained in Section 4.1, however, restricts this broadcasting to $f+1$ parties, significantly reducing communication complexity.
-The protocol replaces the $O(n)$ commit broadcast with a single recommend message.
-We provide a details about classic MVBA protocol challenges in the Appendix B.2.
-Since $f+1$ parties generate verifiable proofs and broadcast their proposals, a party must respond after receiving the first verifiable proof.
-If multiple parties complete the pVCBC protocol, different parties may receive proofs from different parties, hindering agreement.
-Therefore, we introduce the recommend step, where a party broadcasts received proofs to all parties, ensuring at least one proof reaches the majority of the parties ($12f+$ 1), which is necessary to employ ABBA protocol to reach an agreement.
-We use majority and threshold-number alternatively.
-Replacing Commit with Recommend The key idea in pMVBA is to replace the commit step with the recommend step.
-The commit step in classic MVBA protocols involves collecting $n-f$ verifiable proofs and creating an array of length $n$.
-In contrast, the recommend step involves broadcasting verifiable proofs directly, significantly reducing complexity.
-Construction of the Recommend Step The recommend step consists of sending and collecting recommendations.
-This step ensures that at least one verifiable proof reaches the majority of parties.
-The detailed SCR protocol is shown in Algorithm 2:
-  • Initialization: A party initializes an empty set $\Sigma$.
-  • Broadcast Recommendation: A party creates a recommendation message and broadcasts it to all parties.
-  • Recommendation Collection: Upon receiving a recommendation, a party verifies the threshold-sign and instance. If valid, the party adds the proposal and $\sigma$ to its set $\Sigma$.
-  • A party waits for $n-f$ recommendation messages before proceeding.
-1 Local variables initialization:
-2 $\Sigma\leftarrow\{\}$
-3
-4 upon $SCR\langle ID,proposal,\sigma\rangle$ invocation do
-[…]
-6       wait until $|\Sigma|$ = $n-f$
-7       return $\Sigma$
-8
-9
-10
-11upon receiving $\langle recommendation,ID,proposal,\sigma\rangle$ from party $p_{j}$ for the first time do
-12       if  $ID.instance=instance$wedge $Verify_{id}\langle proposal,\sigma\rangle=true$  then
-13             $\Sigma\leftarrow{\langle proposal,\sigma\rangle\cup\Sigma}$
-14
-15
-16
-Algorithm 2 Recommend Protocol for all parties
-Detailed pMVBA Protocol We now present the full details of the pMVBA protocol, inspired by the Cachin-MVBA protocol [13] but with fewer broadcasts.
-The protocol replaces the $O(n)$ commit broadcast with one recommend message and runs an expected constant number of ABBA instances to reach an agreement on a party’s proposal.
-The formal description of the pMVBA protocol is shown in Algorithm 3. It consists of the following steps:
-  • Party/Committee Selection: When a new instance starts, a party invokes the CS protocol to select parties. (See Algorithm 1 lines 1-10 and Algorithm 3 line 10)
-  • pVCBC Broadcast: A selected party broadcasts its proposal using the pVCBC protocol and generates verifiable proofs of their broadcasts. (See Algorithm 3 line 13 and Algorithm 4 lines 1-16.)
-  • Propose and Recommend: Upon completing the pVCBC protocol, a party broadcasts (as a propose) its verifiable proof.
-When a party receives a verifiable proof, it broadcasts the proof as a recommendation and a party waits for $n-f$ recommendation messages.
-(See Algorithm 3 lines 14-22 and the Algorithm 2)
-  • Permutation: When a party receives $n-f$ recommendations, it generates a random order of the selected parties using the cryptographic coin-tossing scheme.
-(See Algorithm 3 line 22 and Algorithm 4 lines 19-29 )
-  • Sequential-ABBA: For the agreement phase, a party runs the ABBA protocol for the selected parties in the determined order, stopping once an agreement is reached.
-(See Algorithm 3 lines 32-58 and Algorithm 5 lines 1-28).
+18    if $id\in RecommendationList$ then
+19
+20      $u_{index}\leftarrow 1$
+21       $proposal,\rho>\leftarrow RecommendationList[id]$
+22       $m_{index}\leftarrow<proposal,\rho>$
+23       send $\langle ID,VOTE,party_{id},u_{index},m_{index}\rangle$ to all parties
+24
+25       else
+26         send $\langle ID,VOTE,party_{id},u_{index},\perp\rangle$ to all parties
+27
+28
+29         wait until $|\Sigma|$ =nf=n-f
+30
+31
+32        if $u_{index}=1$ then
+33           $v\leftarrow\langle 1,<proposal,\rho>\rangle$
+34
+35           else
+36             $v\leftarrow\langle 0,\perp\rangle$
+37
+38             $\langle bit,\rho\rangle\leftarrow ABBA\langle v\rangle$ biased towards 11
+39
+40
+41            if $proposal=\perp$ then
+42               use $\rho$ to complete the verifiable authenticated broadcast and deliver the proposal. See Appendix A.1
+43               return proposal
+44
+45
+46              upon receiving $\langle ID,VOTE,p_{k},u_{index^{\prime}},m\rangle$ for the first time from party $p_{k}$ do
+47                 if $u_{index^{\prime}}=1$ then
+48                   $u_{index}\leftarrow 1$
+49                   $m_{index}\leftarrow m$
+50
+51
+52                  $\Sigma\leftarrow\Sigma\cup 1$
+53
+54
+Algorithm 4 Protocol for party $p_{i}$
+4.7 Integration of Subprotocols
+The pMVBA protocol achieves an agreement in a Byzantine environment through a sequence of interconnected steps.
+The process begins with the Committee Selection (CS) protocol, where each party dynamically and randomly selects committee members (Algorithm 5, line 10).
+Following this, the selected parties broadcast their proposals using the Prioritized Verifiable Consistent Broadcast (pVCBC) protocol, ensuring that their broadcasts are verifiable (Algorithm 5, line 13).
+After the proposals are broadcast, the parties proceed with the Propose-Recommend phase, where verifiable proofs are shared, and recommendations are collected (Algorithm 5, lines 23-27).
+Upon receiving the necessary recommendations, a random order of the selected committee members is generated using a cryptographic coin-tossing scheme (Algorithm 5, line 18).
+Finally, the Sequential-ABBA protocol is executed, where the agreement process is run in the determined order until an agreement is reached on a proposal (Algorithm 5, lines 20-21).
+Each of these steps ensures the protocol operates efficiently and securely, even in the presence of Byzantine faults.
 1
-2 Local variables initialization:
+2Local variables initialization:
 3   $INSTANCE\leftarrow 1$
-4      $PParties\leftarrow\bot$
-5      $\sigma\leftarrow\bot$
+4   $SelectedParties\leftarrow\bot$
+5   $\rho(threshold-signature)\leftarrow\bot$
 6   $Proposal\leftarrow\bot$
-7      $W\leftarrow\bot$
-8      $\Sigma\leftarrow\bot$
+7   $RecommendationList\leftarrow\bot$
+8   $PermutationList\leftarrow\bot$
 9
@@
 12     $ID\leftarrow{\langle INSTANCE,id\rangle}$
-13           $PParties\leftarrow CS\langle ID,instance\rangle$
+13     $SelectedParties\leftarrow CS\langle ID,INSTANCE\rangle$
 14
-15          if $p_{i}\in PParties$ then
-16                $proposal\leftarrow request$
-17                $\sigma\leftarrow pVCBC\langle ID,requests\rangle$
-[…]
+15    if $p_{i}\in SelectedParties$ then
+16       $Proposal\leftarrow requests$
+17       $\rho\leftarrow pVCBC\langle ID,Proposal\rangle$
+18       $\textbf{send}\langle PROPOSE,ID,\rho,Proposal\rangle$ to all parties
 19
-20                wait for a propose or a recommend message.
-21               upon receiving $\langle\textit{propose},ID,\sigma,proposal\rangle$ for the first time do
-22                     $W\leftarrow SCR\langle ID,proposal,\sigma\rangle$
-23
-24
-25                    upon receiving $\langle\textit{recommendation},ID,proposal,\sigma\rangle$ for the first time do
-26                          if no $recommendation$ is sent yet  then
-27                               $W\leftarrow SCR\langle ID,proposal,\sigma\rangle$
-28
-29
-30
+20       wait until $|RecommendationList|\geq 1$
+21
+22
 23      Permutation:
-32                               $\Sigma\leftarrow Permutation\langle ID,PParties\rangle$
-33
+24       $PermutationList\leftarrow RandomOrder\langle ID,SelectedParties\rangle$
+25
 26      Sequential-ABBA:
-35                               $proposal\leftarrow Sequential-ABBA\langle\Sigma,W\rangle$
+27       $Proposal\leftarrow Sequential$-$ABBA\langle PermutationList,RecommendationList\rangle$
 28       $decide\langle Proposal\rangle$
 29       $INSTANCE\leftarrow INSTANCE+1$
-38
+30
+31
+32      upon receiving $\langle\textit{PROPOSE},ID_{k},\rho,Proposal\rangle$ for the first time do
+33         $RecommendationList\leftarrow SCR\langle ID,ID_{k},Proposal,\rho\rangle$
+34
+35
+36        upon receiving $\langle\textit{RECOMMEND},ID_{k},Proposal,\rho\rangle$ for the first time do
+37           if no RECOMMENDATIONRECOMMENDATION is sent yet then
+38             $RecommendationList\leftarrow SCR\langle ID,ID_{k},Proposal,\rho\rangle$
 39
+40
+41
 Algorithm 5 Prioritized-MVBA: Protocol for party $p_{i}$
@@
 In this Section, we conduct extensive evaluation and analysis to answer:
-•  What are the formal proofs supporting the correctness of the Prioritized-MVBA protocol? (§$\S$ 5.1)
-•  How does the Prioritized-MVBA protocol perform in terms of message complexity, communication complexity, and running time? (§$\S$ 5.2)
-•  What are the observed behaviors and outcomes of the protocol under different network conditions and adversarial strategies? (§$\S$ 5.3)
-•  How does the performance of the Prioritized-MVBA protocol compare to other MVBA protocols in terms of key metrics? (§$\S$ 5.4)
+• What are the formal proofs supporting the correctness of the pMVBA protocol? ($\S$ 5.1)
+• How does the pMVBA protocol perform in terms of message complexity, communication complexity, and running time? ($\S$ 5.2)
+• What are the behaviors and outcomes of the protocol under different network conditions and adversarial strategies? ($\S$ 5.3)
+• How does the performance of the pMVBA protocol compare to other MVBA protocols in terms of key metrics? ($\S$ 5.4)
+• How does pMVBA’s performance compare to atomic broadcast protocols in communication complexity? ($\S$ B.1)
 5.1 Proof of the Prioritized-MVBA Protocol
@@
 For the ABBA protocol to output 1, at least one honest party must input 1.
-We substantiate our protocol’s integrity by demonstrating that at least $12f+$ 1 parties receive a verifiable proof of a party’s proposal, thus ensuring the necessary input for the ABBA instance.
+We substantiate our protocol’s integrity by demonstrating that at least $2f+1$ parties receive a verifiable proof of a party’s proposal, thus ensuring the necessary input for the ABBA instance.
 Lemma 5.1.
-At least one party’s proposal reaches $12f+$ 1 parties.
+At least one party’s proposal reaches $2f+1$ parties.
 Proof.
-Since the selected parties can be byzantine and the adversary can schedule the message delivery to delay the agreement, we have considered the following scenarios:
+Since both the selected and non-selected parties can be non-responsive (Byzantine or system failure), and the adversary can coordinate with the Byzantine parties to schedule the message delivery in a way that slows down the agreement process.
+To prove the lemma, we parameterized the number of responsive parties among the selected parties and the total number of responsive parties among both selected and non-selected parties.
+Below is the parameterized statement.
+  • Among the $f+1$ selected parties, $t$ parties are responsive where $1\leq t\leq f+1$ , and among total $n$ parties, $m$ number of parties are responsive, where $2f+1\leq m\leq n$.
+We provide three scenarios below, which are special cases of the above statement.
+We utilize the special cases to prove the parameterized statement.
+The special cases are three edge cases where the maximum number of selected and the non-selected parties can be responsive and non-responsive (case 1 and 2)
     1. 1.
-    Among $\langle f+1\rangle$ selected parties, $f$ parties are non-responsive.
+      Among $f+1$ selected parties, $f$ parties are non-responsive. (The maximum number of selected parties can be non-responsive)
     2. 2.
-    Selected $\langle f+1\rangle$ parties are responsive, but other $f$ non-selected parties are non-responsive.
+      Selected $f+1$ parties are responsive, but other $f$ non-selected parties are non-responsive. (The maximum number of non-selected parties can be non-responsive)
     3. 3.
-    Every party is responsive, including the selected $\langle f+1\rangle$ parties.
-  4. 4.
-    Selected $t<\langle f+1\rangle$ parties are responsive, and total $m$ parties responsive, where $\langle 2f+1\rangle\leq m\leq n$.
+      Every party is responsive, including the selected $f+1$ parties. (The maximum number of responsive parties)
+    We will first prove that the above three scenarios are a special case of the parameterized statement. These three cases are the lower and upper bound for $t$ and $m$.
+    1. 1.
+      For case 1, $t=1$ (selected responsive parties) and $m=t+2f=2f+1$ (since $f$ selected parties are non-responsive, the total number of responsive parties is $2f+1$ that includes one responsive selected party).
+So it adheres requirement of the parameterized equation.
+    2. 2.
+      For case 2, $t=f+1$ and m=t+f=f+1+f(f+1selectedresponsivepartyandfnonseletedresponsiveparty)=2f+1m=t+f=f+1+f(f+1selectedresponsivepartyandfnon-seletedresponsiveparty)=2f+1.
+So it also adheres requirement of the parameterized equation.
+    3. 3.
+      For case 3, $t=f+1$ and $m=t+2f=3f+1=n$. Therefore, it also adheres requirement of the parameterized equation.
 We prove that in every scenario, at least one party’s proposal reaches $\langle 2f+1\rangle$ parties.
+Since we have proved that cases (1), (2), and (3) are the special cases of a parameterized statement, we utilize these cases to prove that even in the scenario of parameterized statement, at least one proposal reaches $2f+1$ parties.
   1. 1.
-    Since among $\langle f+1\rangle$ selected parties, $f$ parties are non-responsive, only one party completes the $pVCBC$ protocol and broadcasts the verifiable proof.
-If any party receives a verifiable proof, then the verifiable proof is from the responsive selected party.
-Since every party receives the verifiable proof for the same party’s proposal, the proposal reaches $\langle 2f+1\rangle$ parties.
+    The proof proceeds by contradiction.
+Assume that among the $f+1$ selected parties, $f$ number of parties are nonresponsive, meaning only one party completes the $pVCBC$ protocol and broadcasts the verifiable proof.
+Under the assumption that no proposal reaches more than $2f$ parties, the total number of recommendations would be:
+    $(1)(2f)(\frac{2f+1}{1})$
+    On average, a selected party’s proposal is recommended by $\frac{2f+1}{1}$ parties. Simplifying this:
+[…]
+    Thus, the total number of recommendations is $(2f)(2f+1)$, which is less than the required $(2f+1)(2f+1)$ recommendations.
+    This leads to a contradiction, as the protocol necessitates a sufficient number of recommendation messages to guarantee progress.
+Given that honest parties are obligated to provide the necessary number of recommendations, it follows that the proposal must have been received by a minimum of $2f+1$ parties.
+Consequently, the assumption that no proposal reaches more than $2f$ parties is incorrect, and at least one selected party’s proposal reaches $2f+1$ parties, thereby enabling the protocol to function as intended.
   2. 2.
-    The $\langle f+1\rangle$ selected parties are responsive and complete the $pVCBC$ protocol.
-Each selected party broadcasts the verifiable proof, and $\langle 2f+1\rangle$ parties receive the verifiable proof ( another $f$ number of parties are non-responsive).
-Any party receives a verifiable proof, recommends the received verifiable proof, and waits for $\langle 2f+1\rangle$ recommendations.
-If a party receives $\langle 2f+1\rangle$ recommendations, then these recommendations include all $\langle f+1\rangle$ parties’ verifiable proofs because among the $\langle 2f+1\rangle$ received recommendations, $\langle f+1\rangle$ number of recommendations are from the selected parties.
-So, every proposal reaches $\langle 2f+1\rangle$ parties.
+    We will prove by contradiction. Assume that no proposal reaches more than $2f$ parties.
+    Given the assumption that a minimum of $2f+1$ parties are responsive and that $f+1$ of these are chosen to propose, it follows that every party must receive a proposal during the proposepropose step.
+Furthermore, no party commences the Sequential-ABBA protocol until it has received a minimum of $2f+1$ recommendations and has secured the random order necessary to advance with Sequential-ABBA.
+This indicates that a total of $(2f+1)(2f+1)$ recommendation messages must be exchanged.
+    Assuming that no proposal is recommended to more than $2f$ parties and that $f+1$ parties submit their requests, the total number of recommendations would be:
+    $(f+1)(2f)(\frac{2f+1}{f+1})$
+    On average, a proposal from a selected party would receive recommendations from $\frac{2f+1}{f+1}$ parties. Simplifying the expression:
+[…]
+    Therefore, the total number of recommendations is $(2f)(2f+1)$, which is lesser to the necessary $(2f+1)(2f+1)$ recommendation messages.
+    This results in a contradiction, as the protocol necessitates the transmission of sufficient recommendation messages for progress to occur.
+Honest parties are required to ensure the delivery of the necessary number of recommendations; thus, at least one proposal must be received by $2f+1$ parties.
+Therefore, our assumption that no proposal reaches more than $2f$ parties is incorrect.
+Consequently, at least one party’s proposal successfully reaches $2f+1$ parties, thereby ensuring the protocol can advance as necessary.
   3. 3.
-    The proof is by contradiction.
-Let no proposal reaches more than $2f$ parties.
-Since we assume every party is responsive, every party receives a proposal in $propose$ step.
-We also know that no party participate in $Sequential-ABBA$ protocol until the party receives $\langle 2f+1\rangle$ recommendations and get the order to Sequential-ABBA.
-There must be a $\langle 3f+1\rangle*\langle 2f+1\rangle$ recommendation messages.
-If no proposal can be recommended to more than $2f$ parties, the total number of recommendations is $\langle 3f+1\rangle*\langle 2f\rangle<\langle 2f+1\rangle*\langle 2f+1\rangle$ (Though a proposal can be recommended by more than one party we assume that every party recommends to the same $2f$ parties otherwise it would fulfil the requirement of $\langle 2f+1\rangle$ proposals).
-But, the honest parties must send enough recommendation messages to ensure the protocol’s progress and the adversary delivers the messages eventually.
-Therefore, at least one party’s proposal reaches $\langle 2f+1\rangle$ parties, a contradiction.
-  4. 4.
-    The proof is by contradiction.
-Let no proposal reaches to more than $2f$ parties.
-If $1\leq t\leq\langle f+1\rangle$ parties distribute their verifiable proof to $\langle 2f+1\rangle\leq m<\langle 3f+1\rangle$ parties and no proposal reaches more than $2f$ parties, then there must be no more than $m*2f$ recommendations.
-However, $m$ parties must receive $m$*$\langle 2f+1\rangle$ recommendations greater than $m*2f$, a contradiction.
-This lemma ensures that if $12f+$ 1 parties receive a verifiable proof, at least one honest party will input 1 to the ABBA instance, ensuring the protocol reaches an agreement on 1.
+    The proof proceeds by contradiction. Assume that no proposal reaches more than $2f$ parties.
+    Since we assume that all parties are responsive, each party receives a proposal during the proposepropose step.
+Additionally, no party engages in the Sequential-ABBA protocol until it has received a minimum of $2f+1$ recommendations and acquired the random order required to execute the Sequential-ABBA protocol.
+Thus, the protocol requires a total of $(3f+1)(2f+1)$ recommendation messages.
+    If no proposal is recommended to more than $2f$ parties and $f+1$ parties propose their requests, then the total number of recommendations must not exceed $(f+1)(2f)$.
+On average, a proposal from a selected party can receive recommendations from $\frac{3f+1}{f+1}$ parties.
+Therefore, the total number of recommendations does not exceed:
+    $\begin{split}(f+1)(2f)(\frac{3f+1}{f+1})\end{split}$ (3)
+    Simplifying:
+[…]
+    Therefore, the total number of recommendations can be expressed as $(2f)(3f+1)$. The necessary quantity of recommendations is $(3f+1)(2f+1)$.
+    While a proposal can be recommended by more than one party, if it is assumed that all recommending parties recommend the same set of $2f$ parties, the condition of receiving $2f+1$ recommendations remains unfulfilled.
+Consequently, for the protocol to advance, it is essential that the honest parties broadcast a sufficient number of recommendation messages, which the adversary will ultimately deliver.
+    Therefore, a minimum of one proposal must be communicated to $2f+1$ parties.
+This contradicts the initial assumption; therefore, at least one party’s proposal has to reach $2f+1$ parties, thereby ensuring the protocol’s advancement.
+We prove that in every scenario, at least one party’s proposal has been shared to $\langle 2f+1\rangle$ parties.
+Having shown that cases (1), (2), and (3) are particular instances of the general parameterized scenario, the proof of these specific cases allow us to extend the result to the parameterized scenario.
+In the three scenarios analyzed, it is evident that with $m$ responsive parties, the total number of recommendations are $m(2f+1)$.
+If $t$ selected parties are responsive, the total number of recommendations can be represented as $t(2f)(\frac{m}{t})$.
+In this context, $(t)(2f)$ indicates that the proposals from each of the $t$ selected parties can reach a maximum of $2f$ parties.
+Meanwhile, $\frac{m}{t}$ denotes the average number of parties that each selected party’s proposal can reach.
+Therefore, we establish the subsequent relationship:
+[…]
+Simplifying further:
+[…]
+This inequality confirms that the number of recommendations from the selected parties is less than the required total number of recommendations, proving that at least one party’s proposal must reach $2f+1$ parties, ensuring protocol progress.
+This lemma ensures that if $2f+1$ number of parties receive verifiable proof, at least one honest party will input 1 to the ABBA instance, ensuring the protocol reaches an agreement on 1.
 Lemma 5.2.
@@
 Proof.
-The proof works by counting the total number $A$ of $vote=0$ messages (line 44 of Algorithm 4) that are generated by honest parties (over all 00 vote in the agreement loop).
-Since every honest party has received a verifiable proof from one party in the $propose$ broadcast step (line 18 of Algorithm 3), the party generates $vote=0$ messages for at most $f$ number of selected parties.
-Thus $A\leq(f)(n-f)$ (all $(n-f)$ votes are zero).
-For the asynchronous binary byzantine agreement protocol to decide 00 for a particular party’s proposal and to cause one more iteration of the loop, at least $\langle f+1\rangle$ honest parties must propose 00 for the binary agreement (line 46 of Algorithm 4).
-Since honest parties only propose 00 if they have received $\langle 2f+1\rangle$ $vote=0$ messages, there must be at least $\langle f+1\rangle$ honest parties who have generated a $vote=0$ message in this iteration.
-Let $R$ denote the number of iterations of the loop where the binary agreement protocol decides 00.
-From the preceding argument, we have $A\geq R(f+1)$.
-Combining these two bounds on $A$, we obtain $R(f+1)\leq f*(n-f)$.
-$\begin{split}R\leq\frac{(f)(n-f)}{f+1}\\ R\leq\frac{(f)(2f+1)}{f+1}\\ R\leq\frac{(f)(f+1+f)}{f+1}\\ R\leq f+\frac{f^{2}}{f+1}\newline \end{split}$ (1)
-However, there are only $\langle f+1\rangle$ number of selected parties’ proposals, and from Lemma 5.1, at least one verifiable proof reaches $\langle f+1\rangle$ honest parties.
-Therefore, $R<\langle f+1\rangle$ and the adversary can cause at most $R+1$ or $\langle f+1\rangle$ number of iterations.
+Since only $f+1$ parties are selected, the number of iterations is bounded above by $f+1$.
+By Lemma 5.1, at least one verifiable proof is delivered to $f+1$ honest parties, which guarantees that the protocol reaches agreement on at least one party’s proposal.
+We therefore analyze the worst case in which the protocol requires exactly $f+1$ iterations.
+If the ABBA execution order of the selected parties is known in advance, an adaptive adversary can schedule message deliveries so that only the proposal of the last party in that order attains recommendations from $2f+1$ number of parties.
+Consequently, for all earlier selected parties, less than the threshold number of parties cast $vote=1$ and the instance decides 0; for the final selected party, at least $f+1$ honest parties cast $vote=1$ and the instance decides $1$.
+.
 ∎
@@
 Proof.
-The binary agreement protocol biased towards $1$ decides 00 if the honest parties input 00.
-From the Sequential-ABBA protocol, each party must receive $\langle n-f\rangle$ $vote=0$ messages in order to input 00 to an ABBA instance and to decide 00.
+The binary agreement protocol biased towards $1$ decides 0 if the honest parties input 0.
+From the Sequential-ABBA protocol, each party must receive $\langle n-f\rangle$ $vote=0$ messages in order to input 0 to an ABBA instance and to decide 0.
 However, from Lemma 5.1 at least $\langle f+1\rangle$ honest parties receive at least one verifiable proof from a selected party $p$, then it is not the case for the party $p\in\overline{A}$.
 This proves the first claim.
-To prove the second claim, let $A$ denote the total number of 00-votes cast by honest parties.
-Since an honest party votes 00 for at most $f$ number of parties proposals, we have $A\leq f(n-f)$.
-On the other hand, a party can cast $vote=0$ for $f+1-|\overline{A}|$ parties proposals, and at least $\langle f+1\rangle$ honest parties cast $vote=0$ to decide 00 then $A\geq(f+1$ $|\overline{A}|)(f+1)$.
-These bounds on $A$ are the same as in Lemma 5.2 with $R=f+1$-$|\overline{A}|$.
-Using the same argument, it follows that $|\overline{A}|\geq 1$, since $R$ is at most $f$ [^1] To prove the Lemma 5.2 and Lemma 5.3, we have used the same technique of Lemma 7 and Lemma 9 from [13].
+To prove the second claim, we refer to the two special cases of Lemma 5.1.
+In case $1$, one selected party is responsive and every party receives the verifiable proof of that selected party; therefore, $|\overline{A}|=1$.
+For case 2, every party receives a recommendation for each selected party.
+Therefore, each selected party’s verifiable proof reaches $2f+1$ parties, thus $|\overline{A}|=f+1$.
+Since Lemma 5.1 proves that at least one party’s proposal reaches $2f+1$ parties, it generalizes the claim $|\overline{A}|\geq 1$.
+The third claim follows now because $|\overline{A}|$ is at least a constant fraction of $f+1$ and thus, there is a constant $\beta>1$ such that $Pr[\Pi(i))\notin\overline{A}]\leq\frac{1}{\beta}$ for all $1\leq i\leq t$.
+Since the probability of the $t$ first elements of $\Pi$ jointly satisfying the condition is no larger than for $t$ independently and uniformly chosen values, we obtain $Pr[\Pi[1]\notin\overline{A}\wedge\Pi[2]\notin\overline{A}\wedge....\wedge\Pi[t]\notin\overline{A}]\leq\beta^{-t}$  [^1] To prove the part of the Lemma 5.3, we have used the same technique of Lemma 9 from [SECURE03].
 We also refer interested readers to Lemma 10 of the same paper for the pseudorandom generation..
-The third claim follows now because $|\overline{A}|$ is at least a constant fraction of $f+1$ and thus, there is a constant $\beta>1$ such that $Pr[\Pi(i))\notin\overline{A}]\leq\frac{1}{\beta}$ for all $1\leq i\leq t$.
-Since the probability of the $t$ first elements of $\Pi$ jointly satisfying the condition is no larger than for $t$ independently and uniformly chosen values, we obtain $Pr[\Pi[1]\notin\overline{A}\wedge\Pi[2]\notin\overline{A}\wedge....\wedge\Pi[t]\notin\overline{A}]\leq\beta^{-t}$.
 Theorem 5.4.
@@
 Agreement: If an honest party outputs a value $v$, then every honest party outputs the value $v$.
-From Lemma 5.1, at least one proposal reaches $12f+$ 1 parties, and the ABBA protocol reaches an agreement on $1$ if $f+1$ honest parties input $1$.
-The agreement property of the ABBA protocol ensures all honest parties output $1$ for that ABBA instance and receive the same value $v$, which satisfies $threshold-validate\langle v,\sigma\rangle=true$.
+From Lemma 5.1, at least one proposal reaches $2f+1$ parties, and the ABBA protocol reaches an agreement on $1$ if $f+1$ honest parties input $1$.
+The agreement property of the ABBA protocol ensures all honest parties output $1$ for that ABBA instance and receive the same value $v$, which satisfies $threshold-validate\langle v,\rho\rangle=true$.
 Liveness: If honest parties participate and deliver messages, all honest parties decide.
-From Lemma 5.1, at least one proposal reaches $12f+$ 1 parties, ensuring the ABBA protocol decides $1$.
+From Lemma 5.1, at least one proposal reaches $2f+1$ parties, ensuring the ABBA protocol decides $1$.
 Lemma 5.3 confirms the protocol reaches an agreement after a constant expected number of iterations.
-External-validity: If an honest party terminates and decides on a value $v$, then $externally-valid\langle v,\sigma\rangle=true$.
-The validity of the ABBA protocol ensures at least one honest party inputs $1$ to the ABBA instance, meaning the honest party received a valid $threshold-signature$ $\sigma$ for $v$.
+External-validity: If an honest party terminates and decides on a value $v$, then $externally-valid\langle v,\rho\rangle=true$.
+The validity of the ABBA protocol ensures at least one honest party inputs $1$ to the ABBA instance, meaning the honest party received a valid thresholdsignaturethreshold-signature $\rho$ for $v$.
 Integrity: If honest parties decide on a value $v$, then $v$ was proposed by a party.
-The ABBA protocol returns $1$ if at least one honest party inputs $1$, which requires a valid $threshold-signature$ $\sigma$.
-Honest parties reply with $sign-shares$ only if a value $v$ is proposed, ensuring $v$ was proposed by a party.
+The ABBA protocol returns $1$ if at least one honest party inputs $1$, which requires a valid thresholdsignaturethreshold-signature $\rho$.
+Honest parties reply with signsharessign-shares only if a value $v$ is proposed, ensuring $v$ was proposed by a party.
 ∎
@@
 Each sub-protocol and step, except for Sequential-ABBA, has a constant running time.
-The running time of the proposed protocol is dominated by the Sequential-ABBA sub-protocol, which has an expected constant number of asynchronous rounds.
+The running time of the proposed protocol is dominated by the Sequential-ABBA sub-protocol.
+The Sequential-ABBA sub-protocol runs the ABBA protocol biased towards $1$ in an expected constant number of times (Algorithm 4 takes a permutation list of the selected parties as an argument and iterates over that list).
+Since we take a permutation of the selected parties (see Algorithm 2 line 19-24), the probability of choosing a selected party is an expected constant number.
+The ABBA protocol also has an expected constant number of asynchronous rounds.
+Therefore, the running time of the protocol is an expected constant number.
 Message Complexity:
 In all sub-protocols and steps, except for pVCBC and propose steps, each party communicates with all other parties.
-Thus, the message complexity is $O(n^{2})$.
+Every party transmit $O(1)$ information to all other parties (See line 8 of Algorithm 1, line 6 of Algorithm 3, lines 14 of Algorithm 5, lines 3, 14, and 22 of Algorithm 2, lines 16 and 18 of Algorithm 4.
+Each of the broadcasts sends $O(1)$ information).
+Since $n$ parties send $O(1)$ information to the $n$ parties, the message complexity is $O(n^{2})$.
 The expected message complexity of the Sequential-ABBA protocol is also $O(n^{2})$.
@@
 The communication complexity of each sub-protocol and step is $O(n^{2}(l+\lambda))$, where $l$ is the bit length of input values and $\lambda$ is the bit length of the security parameter.
+To calculate the communication complexity we use the same approach as message complexity.
+We observe that in no step a party transmit $O(n)$ information.
+Thus, the communication complexity is same as message complexity only includes the bit length of the input values and the bit length of the security parameters.
 The expected communication complexity of the Sequential-ABBA protocol is also $O(n^{2}(l+\lambda))$.
-In conclusion, the Prioritized-MVBA protocol achieves optimal efficiency and correctness by reducing the number of broadcasts and ensuring all security properties of a multi-valued Byzantine agreement protocol.
-The detailed analysis confirms its resilience, integrity, liveness, and external validity, making it suitable for robust decentralized infrastructures.
+Worst-Case Analysis:
+Table 2 presents the analysis of worst-case rounds for each base protocol. The results show that the worst-case rounds are consistent in terms of $f$ across all the base protocols.
 5.3 Case Study
@@
 The main challenge is to provide sufficient information to the parties to maintain the protocol’s progress while removing the costly elements.
-It is crucial to ensure that at least one selected party’s proposal reaches at least $12f+$ 1 parties.
+It is crucial to ensure that at least one selected party’s proposal reaches at least $2f+1$ parties.
 This is achieved through the pVCBC step and the propose-recommend step.
@@
 $p_{1}$’s proposal is received by every party, but node $p_{2}$’s proposal is received by only one party.
+Figure 4: Message flow without the presence of faulty nodes.
 (a) One selected node is non-Byzantine ($p_{1}$), and the adversary delivers the messages from this node to every other node.
@@
 The red node represents either the node is Byzantine or non-responding or the network is not delivering the messages from that node.
+Figure 5: Message flow in the presence of faulty nodes.
 In conclusion, the above figures collectively demonstrate that regardless of the message delivery pattern or the presence of faulty nodes, the protocol consistently achieves agreement, thereby proving its robustness and effectiveness.
@@
 This configuration allows us to explore various message distribution patterns.
-We require at least one party’s proposal to reach $72f+1=$ 7 parties, with a total of $f+1=4$ selected parties.
+We require at least one party’s proposal to reach $2f+1=7$ parties, with a total of $f+1=4$ selected parties.
 We examine the impact of different numbers of faulty parties and message distribution patterns.
@@
 5.4 Analysis of Communication Complexity and Resilience
-This section evaluates our protocol’s performance by comparing its communication complexity with classic MVBA protocols and assessing its resilience, termination, and safety properties against committee-based protocols.
-5.4.1 Comparison of communication complexity with classic protocols.
-Our work focuses on the first polynomial time asynchronous byzantine agreement protocol and removes the expensive sub-component from that protcol.
-We compare our work with the classic MVBA protocol and the traditional MVBA protocols.
-Table 1 highlights these comparisons.
-Our techniques differ from the recent improvements in the MVBA protocol.
-Therefore, we compare our protocol with the classic protocol and the following improvement.
-Our protocol differs from the Cachin-MVBA [13] protocol in terms of communication complexity.
-VABA [23] is a view-based protocol where each view is an instance (a complete execution of the protocol) of the protocol.
-VABA does not guarantee that the parties will reach an agreement at each instance.
-In our protocol, parties can reach an agreement on a valid proposal in every instance.
-We also present a comparison between our proposed protocol and the Dumbo-MVBA [29] protocols.
+This section evaluates our protocol’s performance by comparing its communication complexity with classic MVBA protocol and assessing its resilience, termination, and safety properties against committee-based protocols.
+5.4.1 Comparison of Asynchronous Rounds in MVBA Protocols
+We compare the asynchronous rounds required to reach agreement in our pMVBA protocol with those in classic and traditional MVBA protocols, considering both best-case and worst-case scenarios.
+To determine the worst-case cost, we first calculate the best-case scenario, identify the concrete rounds, and then estimate the remaining rounds as average ABBA rounds, multiplying by $f+1$.
+As shown in Table 2, Dumbo-MVBA requires the most rounds in the best-case scenario.
+In the worst-case, while all protocols require rounds proportional to $f$, VABA struggles to guarantee agreement on a complete output after $f+1$ instances, requiring 13 rounds per instance—twice as many as other protocols.
+Among existing MVBA protocols, pMVBA offers superior performance in communication complexity and asynchronous rounds.
+Although Speeding Dumbo achieves lower asynchronous rounds, it is an ACS protocol and cannot avoid the $O(\lambda n^{3}\log n)$ communication complexity.
+Table 2: Asynchronous rounds of MVBA protocols
+Protocols Rounds (Best case) Rounds (Worst case)
+Cachin et al. [SECURE03] $\boldsymbol{12}$ $\geq 5+7(f+1)$
+VABA [BYZ17] 1313 $\geq 13(f+1)$
+Dumbo-MVBA [BYZ20] 1818 $\geq 11+7(f+1)$
+Our work 1313 $\boldsymbol{6+7(f+1)}$
+5.4.2 Comparison of communication complexity with MVBA protocols.
+Our work focuses on the first polynomial-time asynchronous Byzantine agreement protocol and removes its expensive sub-component.
+We compare our work with the classic and traditional MVBA protocols, as highlighted in Table 3.
+Our techniques differ from recent improvements in MVBA protocols, so we compare our protocol with the classic protocol and the following improvement.
+Our protocol differs from the Cachin-MVBA [SECURE03] in communication complexity.
+VABA [BYZ17] is a view-based protocol where each view is a complete protocol execution.
+VABA does not guarantee agreement in every instance, whereas our protocol ensures parties reach an agreement on a valid proposal in each instance.
+We also present a comparison between our proposed protocol and the Dumbo-MVBA [BYZ20] protocols.
 The Dumbo-MVBA protocol is intended for large inputs.
@@
 We can use the information dispersal technique like Dumbo-MVBA in our protocol to reduce the communication complexity to $O(ln+\lambda n^{2})$, but our protocol does not depend on the erasure code to remove the $O(n^{3})$ term.
-Table 1: Comparison for performance metrics of MVBA protocols
-Protocols Comm. (bit) Word Time Msg. ErasureCode. Instance
-Cachin [13] $O(ln^{2}+\lambda n^{2}+n^{3})$ $O(n^{3})$ $O(1)$ $O(n^{2})$ No 1
-VABA [23] $O(ln^{2}+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ No $\geq 1$
-Dumbo-MVBA [29] $O(ln+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ Yes 1
-Dumbo-MVBA* [29] $O(ln+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ Yes 1
-Our work $O(ln^{2}+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ No 1
+We use a committee to remove the $O(n^{3})$ term it takes only one message and the cryptographic computation.
+The cryptographic computation is similar to other parts of the protocol.
+Table 3: Comparison for performance metrics of MVBA protocols (E. Code stands for Erasure Code and Inst stands for Instance)
+Protocols Comm. (bit) Word Time Msg. E. Code. Inst
+Cachin [SECURE03] $O(ln^{2}+\lambda n^{2}+n^{3})$ $O(n^{3})$ $O(1)$ $O(n^{2})$ No 1
+VABA [BYZ17] $O(ln^{2}+\lambda n^{2})$ $O(n^{2})$ $O(1)$ $O(n^{2})$ No $\geq$ 1
+Dumbo-MVBA [BYZ20] $\boldsymbol{O(ln+\lambda n^{2})}$ $O(n^{2})$ $O(1)$ $O(n^{2})$ Yes 1
+Our work $O(ln^{2}+\lambda n^{2})$ $\boldsymbol{O(n^{2})}$ $\boldsymbol{O(1)}$ $\boldsymbol{O(n^{2})}$ $\boldsymbol{No}$ $\boldsymbol{1}$
 5.4.3 Comparison of Resilience, Termination, and Safety with Committee-Based Protocols
@@
 Table 4 highlights these comparisons.
-COINcidence [19] assumes a trusted setup and violates optimal resilience.
+COINcidence [BYZ19] assumes a trusted setup and violates optimal resilience.
 It also does not guarantee termination and safety with high probability (whp).
-Algorand [20] assumes an untrusted setup, with resilience dependent on network conditions, and does not guarantee termination whp.
-The Dumbo [11] protocol uses a committee-based approach, but its committee-election protocol does not guarantee the selection of an honest party, thus failing to ensure agreement or termination.
+Algorand [BYZ21] assumes an untrusted setup, with resilience dependent on network conditions, and does not guarantee termination whp.
+The Dumbo [FASTERDUMBO] protocol uses a committee-based approach, but its committee-election protocol does not guarantee the selection of an honest party, thus failing to ensure agreement or termination.
 Our protocol achieves optimal resilience and guarantees both termination and safety, as our committee-election process ensures the selection of at least one honest party.
-This guarantees that the protocol can make progress and reach agreement despite adversarial conditions.
+This guarantees that the protocol can make progress and reach an agreement despite adversarial conditions.
 Table 4: Comparison for performance metrics of the committee based protocols
-Protocols n> Termination Safety
-COINcidence [19] 4.5f whp whp
-Algorand [20] * whp w.p. 1
-Dumbo1 [11] 3f whp w.p. 1
-Dumbo2 [11] 3f whp w.p. 1
+Protocols n>> Termination Safety
+COINcidence [BYZ19] 4.5f whp whp
+Algorand [BYZ21] * whp w.p. 1
+Dumbo1 [FASTERDUMBO] 3f whp w.p. 1
+Dumbo2 [FASTERDUMBO] 3f whp w.p. 1
 Our work 3f w.p. 1 w.p. 1
@@
 Though our protocol differs from the atomic broadcast protocol in a number of proposals, we provide a comparison of our protocol with the atomic broadcast protocol in Appendix B.1.
+6 Implementation and Evaluations
+We implemented and evaluated pMVBA protocol using a self-developed gRPC based protocol simulator.
+Along the way, we also implemented and made systematic comparisons with several typical MVBAs, including VABA [BYZ17] and CKPS01 [SECURE03].
+6.1 Performance of pMVBA Protocol
+We extensively evaluated and compared our proposed pMVBA protocol with VABA and Classic MVBA (CKSP01) protocol.
+Implementation details.
+All protocols are written as a single-process Python 3 program.
+The p2p channels among nodes are established using gRPC insecure channel.
+The Python multiprocessing library handles the concurrent tasks.
+Different from [HoneyBadgerBFT] that optimistically skips the verification for threshold signatures, our tests verify all signatures (simulate time to calculate threshold-signatures), which better reflects the actual performance in the worst cases, e.g., with corruptions.
+Test environment.
+The experiments are conducted using gRPC instances with a similar configuration.
+All tests scale up to $10$ instances.
+A transaction in our tests is a string of 1 KB, which approximates the size of 4 basic Bitcoin transactions with one input and two outputs.
+The batch size (also called the system load sometimes) represents the number of transactions proposed by all nodes in a one-shot agreement, and will vary from 1 to $5\times 10^{2}$.
+Result highlights.
+The key information from our experiments is that the pMVBA significantly reduces the latency of Classic MVBA (CKSP01) with the increase of system sizes and attains a throughput $72\%$ higher than CKSP01, almost regardless of the system size, cf.
+Table 5;
+Table 5: Improvements of latency and throughput
+Peak throughput (tx/s) Basic Latency (sec)
+Scale CKSP01 pMVBA CKSP01 pMVBA
+[…]
+[…]
+[…]
+Figure 12: Latency (s) vs Number of Nodes
+Basic latency.
+We set the input batch size as $4$ transactions (i.e., nearly zero) to test the basic latency of pMVBA (and VABA and CKSP01) and show the results in Fig.
+12.
+It is easy to see that the improvement is consistent with the number of nodes.
+Though the basic latency of VABA, CKSP01, and pMVBA is increasing with larger system scales; the improvement made by pMVBA is always significant.
+Throughput at varying scales.
+We then evaluate the throughput of pMVBA at different system scales with varying batch sizes. Along the way, we also compare pMVBA to VABA and CKSP01 and visualize the comparison results in Fig. 13.
+As illustrated in the first three subgraphs in Fig.
+13, pMVBA attains multi-fold improvements relative to CKSP01, disregarding the system scale and batch size.
+First, pMVBA directly inherits the advantages of using the 1-step VCBC protocol, after that, using fewer VCBC instances leads to fewer cryptographic computations and message delay time, which helps it finally achieve a throughput twice that of CKSP01 in the larger batch sizes and larger number of nodes.
+Moreover, carefully analyzing the Fig.
+13, we notice that pMVBA noticeably outperforms both VABA and CKSP01 with larger batch sizes and with the increasing number of nodes.
+The performance improvement is due to the smaller number of computations and messages needed for the pMVBA protocol.
+The peak throughputs of all MVBA protocols are comprehensively shown in the last subgraph of Fig.
+13.
+Although the throughput of all protocols decreases with scaling up, pMVBA always maintains the highest peak throughput among them.
+Figure 13: Throughput of Several MVBA Protocols
+Throughput-latency trade-off.
+Fig.
+14 shows the throughput latency trade-offs in pMVBA, CKSP01, and VABA.
+All the trends are roughly straight-line, indicating that all of them eventually become network-bound.
+pMVBA not only achieves higher peak throughput, but also has a latency that is always smaller than that of VABA and CKSP01 at the same throughput.
+Moreover, the latency gap between pMVBA and VABA is greatly narrowed at some small scales (VABA) and large scales (pMVBA).
+This evidence indicates that our multi-fold improvements greatly expand the applicability of the asynchronous BFT protocol.
+Figure 14: Latency (ts) vs throughput (tx/s) of the MVBA Protocols.
 7 Discussions and Conclusion
@@
 The scheme provides the following algorithms:
-[…]
+  • Key generation algorithm: KeySetup($\{0,1\}^{\lambda},n,f+1)\rightarrow\{UPK,PK,SK\}$.
+Given a security parameter $\lambda$, this algorithm generates a universal public key $UPK$, a vector of public keys $PK:=(pk_{1},pk_{2},\ldots,pk_{n})$, and a vector of secret keys $SK:=(sk_{1},sk_{2},\ldots,sk_{n})$.
 […]
@@
 B.1 Comparison with Atomic Broadcast Protocol
-As discussed earlier, when the inputs of each party are nearly identical, outputting the requests of $n-f$ parties is not a viable solution.
+As discussed earlier, when the inputs of each party are nearly identical, outputting the requests of nfn-f parties is not a viable solution.
 This approach results in higher computational effort without increasing the number of accepted transactions.
@@
 Protocols Communication Complexity
-HB-BFT/BEAT0 [5] $O(ln^{2}+\lambda n^{3}logn)$
-BEAT1/BEAT2 [42] $O(ln^{3}+\lambda n^{3})$
-Dumbo1 [11] $O(ln^{2}+\lambda n^{3}logn)$
-Dumbo2 [11] $O(ln^{2}+\lambda n^{3}logn)$
-Speeding Dumbo [10] $O(ln^{2}+\lambda n^{3}logn)$
+HB-BFT/BEAT0 [HONEYBADGER01] $O(ln^{2}+\lambda n^{3}logn)$
+BEAT1/BEAT2 [DUAN18] $O(ln^{3}+\lambda n^{3})$
+Dumbo1 [FASTERDUMBO] $O(ln^{2}+\lambda n^{3}logn)$
+Dumbo2 [FASTERDUMBO] $O(ln^{2}+\lambda n^{3}logn)$
+Speeding Dumbo [SPEEDINGDUMBO] $O(ln^{2}+\lambda n^{3}logn)$
 Our Work $O(ln^{2}+\lambda n^{2})$
@@
 Upon completion of this step, the party broadcasts both the verifiable proof and the request, providing evidence that the request has been broadcast to every other party.
-Upon receiving verifiable proof from $n-f$ parties, signaling the completion of the VCBC protocol by the threshold number of parties, a party can initiate the ABBA protocol.
+Upon receiving verifiable proof from nfn-f parties, signaling the completion of the VCBC protocol by the threshold number of parties, a party can initiate the ABBA protocol.
 However, there exists the possibility that other parties have not received sufficient verifiable proof, or that the adversary manipulates the distribution of proofs in a manner that prevents the majority of ABBA instances from receiving adequate proof.
-To address this, each party communicates with others by transmitting an $n$-bit array, indicating receipt of verifiable proof from $n-f$ parties.
-Upon receipt of $n-f$ verifiable proofs, a party generates a permutation of the parties and invokes ABBA instances based on the order of permutation, ensuring that the number of ABBA instances remains constant on average.
+To address this, each party communicates with others by transmitting an $n$-bit array, indicating receipt of verifiable proof from nfn-f parties.
+Upon receipt of nfn-f verifiable proofs, a party generates a permutation of the parties and invokes ABBA instances based on the order of permutation, ensuring that the number of ABBA instances remains constant on average.
 Appendix C Deferred protocols
-Construction of the pVCBC
-  • Upon invocation of the pVCBC protocol, a party creates a message using the $ID$ and the requests.
-The party then authenticates the message using its private key and multicasts the message (lines 03-06).
-  • Upon receiving a signed message ($msg_{signed}$) from a party $p_{j}$, a party checks whether the sender is a selected party and whether the message is authentic (signed by the sender).
-If the sender is a selected party and the message is authentic, the party adds its signature share ($sign-share$) to the message, resulting in $\rho$.
-The party then replies with $\rho$ to the sender (lines 13-16).
-  • Upon receiving a signature share $\rho_{k}$ from a party $p_{k}$, a selected party checks the validity of $\rho_{k}$ (ensuring it was added using $p_{k}$’s key).
-If $\rho_{k}$ is valid, the party adds the signature share $\rho_{k}$ to its set $\Sigma$ (lines 09-11).
-  • A selected party waits for $\langle n-f\rangle$ valid signature shares.
-Upon receiving $\langle n-f\rangle$ signature shares, the party combines them to generate a threshold signature $\sigma$ (proof that the party has sent the same request to at least $\langle f+1\rangle$ honest parties) and returns $\sigma$ to the caller (lines 07-08).
 Construction of the Permutation
 Here is the pseudocode for the permutation protocol (see Algorithm 2). Below is a step-by-step description of the protocol:
-  • Upon invocation of the Permutation protocol, a party generates a coin-share $\sigma_{i}$ for the instance and broadcasts $\sigma_{i}$ to every party, then waits for $12f+$ 1 coin-shares (lines 22-24).
+  • Upon invocation of the Permutation protocol, a party generates a coin-share $\sigma_{i}$ for the instance and broadcasts $\sigma_{i}$ to every party, then waits for $2f+1$ coin-shares (lines 22-24).
   • When a party receives a coin-share from another party $p_{k}$ for the first time, it verifies the coin-share (ensuring it is from $p_{k}$) and accumulates the coin-share in the set $\Sigma$.
-The party continues to respond to coin-shares until it has received $12f+$ 1 valid shares (lines 27-29).
-  • Upon receiving $12f+$ 1 valid coin-shares, a party uses its $CToss$ function and the collected coin-shares to generate a permutation of the $n$ parties (lines 24-25).
-Sequential-ABBA
-The purpose of the Sequential-ABBA protocol is to run an agreement loop to agree on a selected party’s proposal.
-We use the asynchronous binary Byzantine agreement protocol biased towards $1$ for each selected party until the parties reach an agreement on one of the party’s proposals.
-Construction of the Sequential-ABBA
-Since the Sequential-ABBA protocol guarantees that the parties will eventually reach an agreement on one of the selected party’s proposals, the protocol runs a loop for the selected parties until it agrees on a valid proposal.
-The pseudocode of the Sequential-ABBA protocol is given in Algorithm 5, and a step-by-step description is provided below:
-  • The Sequential-ABBA protocol takes two arguments: (i) $\Sigma$, the permutation of the selected parties, and (ii) $W$, the list of verifiable proofs the party has received (line 1).
-  • Upon invocation of the protocol, a party declares two variables: (i) $l$, an index number to access the selected parties one by one from the array $\Sigma$, and (ii) $b$, initially set to zero to indicate that the parties have not reached any agreement (lines 2-3).
-  • While loop: Parties are chosen one after another according to the permutation $\Sigma$ of $\{1,\ldots,f+1\}$.
-Let $l$ denote the index of the party selected in the current loop (the selected party $p_{l}$ is called the candidate).
-Each party $p_{i}$ follows these steps for the candidate $p_{l}$ (lines 4-22):
-    • Broadcasts a $vote$ message to all parties containing $u_{l}=1$ if party $p_{i}$ has received $p_{l}$’s proposal and verifiable proof (including the proposal in the vote), and $u_{l}=0$ otherwise.
-    • Waits for $\langle n-f\rangle$ $vote$ messages but does not count votes indicating $u_{l}=1$ unless a valid proposal from $p_{l}$ has been received—either directly or included in the $vote$ message.
-    • Runs a binary validated Byzantine agreement biased towards $1$ (see Algorithm 6) to determine whether $p_{l}$ has properly broadcast a valid proposal.
-Vote $1$ if $p_{i}$ has received a valid proposal from $p_{l}$ and adds the protocol message that completes the verifiable broadcast of $p_{l}$’s proposal to validate this vote.
-Otherwise, if $p_{i}$ has received $n-f$ $vote$ messages containing $u_{l}=0$, then vote 00; no additional information is needed.
-If the agreement loop decides $1$, exit the loop.
-  • Upon reaching an agreement, a party returns the threshold signature of the agreed proposal.
-  • Upon receiving $\langle ID,vote,party,u_{l},m\rangle$, a party checks whether $u_{l}=1$. If it receives $u_{l}=1$, then it assigns $m_{l}=m$.
+The party continues to respond to coin-shares until it has received $2f+1$ valid shares (lines 27-29).
+  • Upon receiving $2f+1$ valid coin-shares, a party uses its $CToss$ function and the collected coin-shares to generate a permutation of the $n$ parties (lines 24-25).
+Appendix D Agreement protocol
+D.1 Asynchronous Binary Byzantine Agreement (ABBA)
+The ABBA protocol allows parties to agree on a single bit $b\in\{0,1\}$ [BYZ10, SECURE05, SIG01].
+We have adopted the ABBA protocol from [SECURE02], as given in Algorithm 6.
+The expected running time of the protocol is $O(1)$, and it completes within $O(k)$ rounds with probability $1-2^{-k}$.
+Since the protocol uses a common coin, the total communication complexity becomes $O(kn^{2})$.
+For more information on how to realize a common coin from a threshold signature scheme, we refer interested readers to the [HONEYBADGER01].
+Construction of the ABBA biased towards 1
+We use the ABBA protocol from [SECURE02].
+We optimize and changed the protocol for biased towards $1$.
+The biases towards $1$ property ensures that if at least one party input $1$ in the pre-process step.
+The pseudocode of the ABBA protocol biased towards 1 is given in Algorithm 6, and a step-by-step description is provided below:
+  • Pre-process step . Generate an $\sigma_{0}$ share on the message and send the pre-process type message to all parties.
+  • Collect $2f+1$ proper pre-processing messages. (see (Algorithm 6)).
+  • Repeat loop: Repeat the following steps 1-4 for rounds round = 1,2,3,…
+    • Pre-Vote step. (see Algorithm 7)
+        If round = 1, $b=1$ if there is a pre-processing vote for $1$ (biased towards 1, taking one vote instead of majority) else $b=0$. (see lines 3-4).
+        If round >> 11, if there is a threshold signature on main-vote message from round-1 then decide and return. (see lines 18-20)
+        Upon receiving main-vote for $0/1$, update $b$ and the justification. (see lines 12-17)
+        b=F(ID,r1)b=F(ID,r-1), all the main-vote are abstain and the justification is the threshold signature of the abstain vote. (see lines 6-7)
+        Produce signature-share on the message (ID, pre-vote, round, b) and send the message of the form pre-vote,round,b,justification, signature-share). (lines 9-11)
+    • Main-vote step. (See Algorithm 8)
+        Collect (2f+1) properly justified round pre-vote messages. (lines 14-19)
+        If there are (2f+1) pre-votes for 0/1, $v=0/1$ and the justification is the threshold-signature of the the sign-shares on pre-vote messages. (lines 5-7)
+        If there are (2f+1) pre-votes for both 0 and $1$, $v=abstain$ and the justification is the two sign-shares from pre-vote 0 and pre-vote 1. (lines 9-10)
+        Produce signature-share on the message (ID, pre-vote, round, v) and send the message of the form (main-vote,round,v,justification, signature-share) to all parties (lines 11-13)
+    • Check for decision. (See Algorithm 9)
+        Collect (2f+1) properly justified main-votes of the round roundround. (line 3)
+        If these is no abstain vote, all main-votes for $b\in\{0,1\}$, then decide the value $b$.
+Produce a threshold signature on the main votes’ sign-shares and send the threshold signatures to all parties and return.
+(lines 4-7)
+        Otherwise, go to Algorithm 9. line (11)
+    • Common Coin. (See Algorithm 10)
+        Generate a coin-share of the coin (ID, round) and send to all parties a message of the form (coin, round,coin-share). (lines 1-4)
+        Collect (2f+1) shares of the coin (ID,round $\sigma_{k}$), and combine these shares to get the value $F(ID,round)\in\{0,1\}$. (lines 5-6)
+1
+2
+3upon ABBA(m) do
+   /* Preprocess Step. */
+4
+5  $\sigma_{0}\leftarrow SigShare_{i}(sk_{i},m_{i})$ see [SECURE02]
+6
+7  send $(pre$-$process,m_{i},\sigma_{i})$ to all parties
+8
+9  wait until at least (nf)(n-f) pre-process messages have been received.
+10
+11  for $round\leftarrow 1,2,3,$... do
+12     Prevote Step : Algorithm 7
+13
+14    Main-vote Step: Algorithm 8
+15
+16    Check For Decision : Algorithm 9
+17
+18    Common Coin: Algorithm 10
+19
+20
+Algorithm 6 ABBA biased towards 1: protocol for party $p_{i}$
+[…]
+2 $justification\leftarrow\perp$
+3 if round = 1 then
+4   $b\leftarrow 1$ if there is any pre-process message with $m=1$ (biased towards 11), otherwise 0.
+5   else
+6     $b\leftarrow F(ID,round-1)$
+7     $justification\leftarrow threshold$-$signature\langle ID,main$-$vote,round-1,abstain\rangle$
+8
+9     wait for nfn-f justified main-vote
+10
+11
+12    $m_{i}\leftarrow(ID,pre$-$vote,round,b)$
+13     $\sigma\leftarrow SigShare_{i}(sk_{i},m_{i})$
+14
+15    send $(pre$-$vote,r,b,justification,\sigma)$ to all parties
+16
+17    upon receiving $\langle main$-$vote,round,v,justification,\sigma\rangle$ for the first time from party $p_{k}$ do
+18       if v = 0 then
+19         $b\leftarrow 0$
+20
+21         else if v = 1 then
+22           $b\leftarrow 1$
+23
+24
+25          $justification\leftarrow threshold$-$signature\langle ID,pre$-$vote,round-1,b\rangle$
+26
+27          upon receiving $\langle b,threshold$-$signature\rangle$ for the first time from party $p_{k}$ do
+28             send(b, threshold-signature) to all parties
+29             return(b, threshold-signature)
+30
+31
+Algorithm 7 ABBA biased towards 1: Pre-vote step
+1
+2
+/* Mainvote Step. */
+3
+[…]
+5 $PV_{0}\leftarrow\{\}$
+6 $PV_{1}\leftarrow\{\}$
+7 wait until $|\Sigma|$ = 2f+1
+8
+9if $|PV_{0}|=2f+1$ or $|PV_{1}|=2f+1$ then
+10   $v\leftarrow 0/1$
+11   $justification\leftarrow CombineShare_{i}(v,{i,\sigma_{i}}_{i\in\Sigma})$
+12
+13   else
+14     $v\leftarrow abstain$
+15     $justification\leftarrow(\sigma_{i}\in PV_{0},\sigma_{j}\in PV_{1})$
+16
+17
+18    $m_{i}\leftarrow(ID,main$-$vote,round,v)$
+19     $\sigma\leftarrow SigShare_{i}(sk_{i},m_{i})$
+20
+21    send $(main$-$vote,round,v,justification,\sigma)$ to all parties
+22
+23    upon receiving $\langle pre$-$vote,round,b,justification,\sigma\rangle$ for the first time from party $p_{k}$ do
+24       if b = 0 then
+25         $PV_{0}\leftarrow PV_{0}+1$
+26
+27         else if b = 1 then
+28           $PV_{1}\leftarrow PV_{1}+1$
+29
+30           $\Sigma\leftarrow\Sigma\cup\{\sigma_{k}\}$
+31
+32
+Algorithm 8 ABBA biased towards 1: Main-Vote step
+1
+2
+[…]
+4 $isAbstain\leftarrow no$
+5 wait until $|\Sigma|$ = 2f+1
+6 if isAbstain = no then
+7   threshold-signature = $CombineShare_{i}(b,{(i,\sigma_{i})}_{i\in\Sigma})$
+8   send(threshold-signature) to all parties
+9   return (b, threshold-signature)
+10
+11   else
+12     go to Algorithm 10
+13
+14
+15    upon receiving $\langle main$-$vote,r,v,justification,\sigma\rangle$ for the first time from party $p_{k}$ do
+16       if v = abstain then
+17         $isAbstain\leftarrow yes$
+18
+19         $b\leftarrow v$
+20         $\Sigma\leftarrow\Sigma\cup\{\sigma_{k}\}$
+21
+22
+Algorithm 9 ABBA biased towards 1: Check for Decision for party $p_{i}$
 1
 […]
-3      $\Sigma\leftarrow\{\}$
-4      $msg\leftarrow createMsg\langle ID,requests\rangle$
-5      multi-cast $\langle msg\rangle$
-6      wait until $|\Sigma|=n-f$
-7      return $\sigma\leftarrow CombineShare_{id}\langle requests,\Sigma\rangle$
-8
+3 $\sigma_{i}\leftarrow CShare(r_{i})$
+4 send $\langle coin,round,\sigma_{i}\rangle$ to all parties
+5 wait until $|\Sigma|$ = 2f+1
+6 $F(ID,round)\in\{0,1\}\leftarrow CToss(r,\Sigma)$
+7
+8upon receiving $\langle coin,round,\sigma_{k}\rangle$ for the first time from party $p_{k}$ do
 9
-10
-[…]
-12           if $VerifyShare_{k}\langle m,(k,\sigma_{k})\rangle$  then
-13                $\Sigma\leftarrow{\sigma_{k}\cup\Sigma}$
-14
-15
-16                upon receiving $\langle msg\rangle$ from the party $p_{j}$ for the first time do
-17                     if $p_{j}\in PParties$  then
-18                          $requests\leftarrow msg.requests$
-19                          $\sigma_{id}\leftarrow SigShare_{id}\langle sk_{id},requests\rangle$
-20                          $reply\langle\sigma_{id}\rangle$
-21
-22
-23
-24
-25
-26                          Local variables initialization:
-27                               $\Sigma\leftarrow\{\}$
-28
-29
-30                               upon $Permutation\langle id,instance\rangle$ invocation do
-31                                    $\sigma_{i}$leftarrow $CShare\langle r_{id}\rangle$
-32                                    $multi-cast\langle SHARE,id,\sigma_{id},instance\rangle$
-33                                    wait until $|\Sigma|=2f+1$
-34
-35                                   return $CToss\langle r_{id},\Sigma\rangle$
-36
-37
-38                                    upon receiving $\langle SHARE,id,\sigma_{k},instance\rangle$ from a party $p_{k}$ for the first time do
-39                                         if $CShareVerify\langle r_{k},k,\sigma_{k}\rangle=true$  then
-40                                              $\Sigma\leftarrow{\sigma_{k}\cup\Sigma}$
-41
-42
-43
-44
-Algorithm 4 Protocol for party $p_{i}$
-1
-[…]
-3      $l\leftarrow 0$
-4      $b\leftarrow 0$
-5      while b=0 do
-6           $\Sigma_{vote}\leftarrow\{\}$
-7           $l\leftarrow l+1$
-8           $party\leftarrow\Sigma\langle l\rangle$
-9           $u_{l}\leftarrow 0$
-10           $proposal_{l}\leftarrow\perp;$ where $11\leq l\leq f+$ 1
+10  $\Sigma\leftarrow\Sigma\cup\{\sigma_{k}\}$
 11
-12          if $l\in W$ then
-13
-14               $u_{l}\leftarrow 1$
-15                $m_{l}\leftarrow W[l]$
-16                multi-cast $\langle ID,vote,party,u_{l}=1,m_{l}\rangle$
-17
-18                else
-19                     multi-cast $\langle ID,vote,party,u_{l}=0,\perp\rangle$
-20
-21
-22                    wait until $|\Sigma_{vote}|$ $=n-f$
-23
-24
-25                    if $u_{l}=1$ then
-26                          $v\leftarrow\langle 1,m_{l}\rangle$
-27
-28                          else
-29                               $v\leftarrow\langle 0,\perp\rangle$
-30
-31                               $\langle b,\sigma\rangle\leftarrow ABBA\langle v\rangle$ biased towards $1$
-32
-33                               return $\langle b,\sigma\rangle$
-34
-35
-36                              upon receiving $\langle ID,vote,party,u_{l},m\rangle$ for the first time from party $p_{k}$ do
-37                                    if $u_{l}=1$  then
-38                                         $u_{l}\leftarrow 1$
-39                                         $m_{l}\leftarrow m$
-40
-41                                        $\Sigma_{vote}\leftarrow\Sigma_{vote}+1$
-42
-43
-Algorithm 5 Protocol for party $p_{i}$
-Appendix D Agreement protocol
-D.1 Asynchronous Binary Byzantine Agreement (ABBA)
-The ABBA protocol allows parties to agree on a single bit $b\in\{0,1\}$ [27, 36, 2].
-We have adopted the ABBA protocol from [3], as given in Algorithm 6.
-The expected running time of the protocol is $O(1)$, and it completes within $O(k)$ rounds with probability $1-2^{-k}$.
-Since the protocol uses a common coin, the total communication complexity becomes $O(kn^{2})$.
-For more information on how to realize a common coin from a threshold signature scheme, we refer interested readers to the paper [5].
-1 upon receiving $b_{input}$, set $est_{0}:=b_{input}$ and proceed as follows in consecutive epochs, with increasing rounds $r$ : do
-2
-3     multi-cast $BVAL_{r}(est_{r})$
-4      bin_valuesr := {}
-5      upon receiving $BVAL_{r}(b)$ messages from f+1 parties do
-6           if $BVAL_{r}(b)$ has not been sent then
-7                multi-cast $BVAL_{r}(b)$
-8
-9
-10
-11               upon receiving $BVAL_{r}(b)$ messages from 2f+1 parties do
-[…]
-13
-14
-15                    while $bin\_values_{r}\neq\phi$ do
-16
-17                         multi-cast $AUX_{r}(v)$, where $v\in bin\_values_{r}$
-18                          wait until at least (n-f) $AUX_{r}$ messages have been received, such that the set of values carried by these messages,$vals$ are a subset of $bin\_values_{r}$.
-19                          $S\leftarrow Coin_{r}.GetCoin()$ see [5, 16]
-20                          if vals = {b} then
-21                               $est_{r+1}:=b$
-22
-23                               if (b = S%2) then
-24                                    output b
-25
-26
-27                                    else
-28                                         $est_{r+1}:=S\%2$
-29
-30
-31                                         continue looping until both a value b is output in some round r, and the value $Coin_{r^{\prime}}=b$ for some round $r^{\prime}>r$.
-32
-Algorithm 6 ABBA: protocol for party $p_{i}$
+12
+Algorithm 10 ABBA biased towards 1: Common Coin for party $p_{i}$